### v1

# Exploring a New Design Paradigm for Omnidirectional MAVs for Minimal Actuation and Internal Force Elimination: Theoretical Framework and Control

Abstract
This paper presents a novel concept for achieving omnidirectionality in a multirotor aerial vehicle (MAV) that uses only 6 inputs and ensures no internal forces at the equilibria. The concept integrates a single actively-tilting propeller along with 3 pendulum-like links, each carrying a propeller, connected by passive universal joints to the main body. We show that this design ensures omnidirectionality while minimizing the internal forces and without resorting to overactuation (i.e., more than 6 inputs). A detailed dynamic model of the multi-link MAV is first developed. Afterwards, the analysis identifies the equilibrium configurations and illustrates that a forced equilibrium exists for every pose of the MAV’s main platform. In order to render this equilibrium asymptotically stable for the closed-loop system, a geometric nonlinear controller is constructed using dynamic feedback linearization and backstepping techniques with the main platform configuration error being the left-trivialized error on $SE(3)$. The stability of the closed-loop system is then investigated by employing standard Lyapunov arguments on the zero dynamics. We conclude by providing numerical simulations validating the proposed approach. They demonstrate the MAV capability to perform decoupled attitude and translational motions under non-zero initial conditions, parametric uncertainty, and actuators noise.
I introduction
Nowadays, Multirotor Aerial Vehicles (MAVs) are widely used in a variety of applications, ranging from agriculture to civil engineering  [1]. Interest in this technology is growing, geared by its potential to perform tasks that require physical interaction with environments inaccessible to other types of mobile vehicles [2, 3]. While conventional quadrotors can efficiently handle some of these tasks, particularly when special orientation maneuvers are not required, certain applications have emerged in which the capability of executing decoupled translational and rotational motions is essential, necessitating the demand for new MAV designs since quadrotor dynamics inherently prohibits this kind of motion. One notable example where the MAV platform needs to maintain constant orientation while translating in space is the contact-based inspection of sloped surfaces  [4, 5].
I-A Research Problems and Related Work
When designing a new MAV, one can easily be convinced that it is desirable to obtain all the angular velocities of the propellers aligned with each other and with the gravity force at any static hovering equilibrium: this avoids internal propeller-generated forces canceling each other at the cost of a higher energy consumption; moreover, it is desirable to keep the number of actuators to the minimum (equal to the task dimension), to reduce the payload (which, in turn, affects the vehicle’s endurance), the cost, the maintenance requirements, etc. Furthemore, for tasks requiring high dexterity like the ones mentioned before, it is desirable for the MAV to be able to statically hover at any pose (in SE(3) in general). Last but not least, the design of a MAV must also allow the controlled stabilization of its hovering equilibria. We label these four desirable properties as the Internal-force Property, the Input-dimension Property, the Equilibria Property, and the Stabilizability Property, respectively.
When restricting the problem to the vertical 2D space, we demonstrated in [6] that a MAV design with all the above four properties exists. Especially, it can be stabilized at any static equilibrium pose in SE(2) with propellers parallel to each other and the gravity vector, and it has only three actuators. The actuators are two brushless motors attached to one propeller each and one servomotor that actively tilts one of the two propellers, while the other one is free to passively rotate and has a center of mass not coincident with the rotation point. However, extending such a result to the 3D space is very far from being trivial and is an open problem. Just to mention two obstacles to such extension: SE(3) is a fundamentally different manifold than SE(2) and the dynamic effects related to single revolute joints in 2D cannot be extended to the case of multiple joints in 3D, see model in Sec. II-B. Moreover, in the 3D space, the propeller drag torques arise and have to be taken into account, see discussion in Sec. III-A.
We will now revise the main MAV designs proposed in the literature and show that none of them has all the four desirable properties described above.
Standard underactuated MAVs like quadrotors and hexarotors tick the Internal-force Property as they have all propellers fixedly oriented with spinning axes parallel to each other; however, because of the fixed attachment between the propellers and the body frame, their only hovering equilibria are characterized by the attitudes where the $z$ axis is aligned with the gravity vector. Thus, they do not tick the Equilibria property.
Aerial vehicles equipped with the ability to hover while maintaining a larger set of orientations, known as fully-actuated MAVs, are surveyed in [7]. These MAVs cannot be stabilized at any pose but at a subset of poses in SE(3) determined by the thrust limits, the design kinematics, and joint limits. Such subset contains any position but only a subset of all the orientations in general. A subclass of this type of MAVs, called omnidirectional MAV [8], can hover at any pose: namely, they have the Equilibria property mentioned above. These aerial vehicles can be divided in two classes.
The first class of designs that have been proposed to achieve full actuation or omnidirectionality are realized by fixedly tilting the propellers in such a way that their spinning axes are not all parallel to each other. As a consequence, all of these designs do not have the Internal-force Property. Examples are [9], a fully actuated hexarotor, and [10, 11], omnidirectional hexarotors. Other omnidirectional fixedly-tilted-propeller designs do not comply with the Input-dimension Property, either; they are, for example, the hepta-rotor presented in [12] or the octarotors proposed in [13, 14].
The second design approach to achieve full-actuation or omnidirectionality while also partially achieving the Internal-force Property for at least some hovering orientations is the addition of servomotors that actively change the relative orientation between the propellers and the frame of the MAV. In [15, 16], the authors present a modification of the conventional quadrotor by actuating the tilt of each propeller separately via 4 servomotors for a total of 8 actuators. Similarly, in [17], the authors used 6 servomotors in addition to the 6 brushless motors that rotate the 6 propellers of an hexarotor. The design in [18] also falls in this category. They developed a MAV with multiple links connected in series by active revolute joints, with each link having two propellers that rotate simultaneously using a motor. In an attempt to limit the number of actuators, some designs have been proposed in which a single servo-motor is used to synchronously tilt all the propellers: 6 propellers are synchronously tilted by the additional actuator in [19], and 8 propellers in [20]. All the aforementioned designs do not comply with the Input-dimension Property, having a number of actuators greater than 6. Furthermore they do not comply with the Internal-force Property at any orientation but only some at some specific ones.
I-B Contributions
The main contribution of this work is to propose the first MAV design, to the best of the authors’ knowledge, having the following properties in SE(3).
  1. 1.
    Input-dimension property. The number of actuators in the proposed MAV design is 6, equal to the number of DoFs of its base’s pose. There are 4 brushless motors, each making one propeller spin; 3 of the propellers are connected to a passive 2-DoF joint; one of the propeller’s attitude is actively controlled by 2 servomotors.
  2. 2.
    Equilibria property. In Proposition 1, we show that static hovering is feasible at any pose in $SE(3)$ of the MAV’s main body; in other words, a forced equilibrium exists for any base’s pose.
  3. 3.
    Internal-force property. Proposition 1 also shows that no internal forces are produced at the equilibrium for any base pose in $SE(3)$.
  4. 4.
    Stabilizability property. We prove in Theorem 1 that it is possible to create a geometric state feedback law that renders the closed-loop system of the MAV asymptotically stable at any base pose in $SE(3)$, as long as configuration and thrust trajectories of the system remain in a certain neighborhood of that base pose. We observe that the dynamics of the passive propeller joints coincide with the zero dynamics, and we prove its asymptotical stability using the Lyapunov theory. The feedback control law is given in Proposition 2.
Eventually, simulation results of the closed-loop system in the presence of non-idealities, such as noise and parameter uncertainties, are presented, in order to provide an acceptable level of confidence in the generalizability of the result beyond the standard model used to formally prove all the properties.
I-C Motivations Behind the Design Idea
The simplest MAV design that one could think of that respects the Input-dimension Property and the Internal-force Property (and potentially the Equilibria Property) is the one with 6 propellers, all free to passively tilt like 3D pendulums. However, the results achieved for the 2D example in [6] showed that a design with 3 passively linked propellers in a pendulum like fashion, in that case, was non-feedback linearizable as the allocation matrix was structurally rank-deficient at the static equilibria. Based on such a preliminary result, we discarded the all-passive-propeller design.
Hence, our design has servomotors that actively reorient one propeller. An alternative could be to reorient actively all the four propellers synchronously by means of the two servomotors and a transmission mechanism. However, the design of such a mechanism to drive the whole propeller joints synchronously is not trivial, would bring mechanical complexity and with it additional payload, cost, and maintenance effort; furthermore, it would not scale well with the size of the vehicle, when the motion needs to be transferred over longer distances. Instead, the proposed design where the servomotors exert their actuation locally at one individual propeller has the advantage of being modular and simple from a mechanical point of view, not posing challenges when increasing the dimensions of the vehicle.
I-D Outline
The rest of the paper is organized as follows. Sec. II describes the dynamic model of the proposed novel MAV conceptual idea, whereas Sec. III identifies the equilibria sets of that model being all the poses in SE(3) and all having the internal force property. Afterwards, we investigate and demonstrate that the system can be made asymptotically stable at all these equilibria by a state feedback in Sec. IV. Numerical simulations where the MAV performs decoupled motions are the subject of Sec. V. We conclude by giving some remarks on the future work in Sec. VI. For ease of exposition, all the formal proofs are collected in a separate section: Sec. VII.
II Proposed Design and Dynamical Model
Figure 1: Conceptual representation of the proposed MAV idea. The origins of the frames $\mathcal{F}_{b}$ and $\mathcal{F}_{p_{2}}$ are attached to CoMs of the base $B_{1}$ and the propeller link $B_{2}$, respectively. The shown configuration is when $\mathcal{F}_{b}$ coincides with $\mathcal{F}_{w}$ and all the frames of the links have the same orientation as $\mathcal{F}_{w}$. The joint $q_{j1}$ and $q_{j2}$ positions are zero when $\mathcal{F}_{p_{j}}$ has the same orientation as $\mathcal{F}_{b}$, $j=\{2,\ldots,5\}$. The link $B_{5}$ is attached to $B_{1}$ through an active joint, unlike the the rest of the joints which are passive.
In this section, we describe the proposed design and derive its dynamical model in state-space representation.
II-A Mechanical Design Idea
The vehicle concept, depicted in Fig. 1, is composed of a main rigid body called the base and denoted with $B_{1}$, and four smaller bodies, referred to as links and denoted with $B_{2},B_{3},B_{4}$ and $B_{5}$. Each link is connected to the base by means of two revolute joints, thoroughly described in the following, and hosts a rigidly attached velocity controlled brushless (BL) motor driving a propeller. We refer as the actuation unit to each sub-system composed by the propeller, the BL motor, the link, and the two revolute joints connecting the link with the base, for a total of four actuation units. The rotation axes of the two revolute joints of each actuation unit are assumed to intersect in one point, called the center of rotation of the corresponding link, which also coincides with the center of the propeller.
The CoM’s of $B_{2}$, $B_{3}$, and $B_{4}$ are located substantially below their centers of rotation. This can be achieved for example by lowering enough the position of the BL motors of these actuation units, which typically are the heaviest parts. This design makes the three links resemble three spatial pendula with the corresponding joints acting as their pivots. The CoM of $B_{5}$ is instead assumed to be coincident with its center of rotation, for a reason that will be clear in the following.
The actuation units are mounted on the base in such a way that their centers of rotation (i.e., the propeller thrust application points) are all placed on the same plane. This is referred to as a coplanar design.
The two revolute joints connecting $B_{5}$ with $B_{1}$ are assumed to be actuated by two servomotors, representing together a universal joint denoted with $J_{15}$, whose configuration is denoted with $q_{5}\in\mathbb{T}^{2}$. The six revolute joints connecting $B_{1}$ with $B_{2}$, $B_{3}$, and $B_{4}$, are passive, representing three universal joints denoted with $J_{12}$, $J_{13}$, and $J_{14}$, respectively, whose configurations are denoted with $(q_{2},q_{3},q_{4})\in\mathbb{T}^{6}$.
For what has been said so far, the generic universal joint $J_{1j}$, with $j=2,\ldots,5$, is then constituted by two consecutive 1-DoF mutually-orthogonal revolute joints whose configurations are represented by $q_{j1}\in\mathbb{S}^{1}$ and $q_{j2}\in\mathbb{S}^{1}$. We assign body-fixed reference frames to $B_{1},B_{2},\ldots,B_{5}$ with origins coincident with the corresponding CoM’s, and we denote them with $\mathcal{F}_{b}$, $\mathcal{F}_{p2}$, $\mathcal{F}_{p3}$, $\mathcal{F}_{p4}$ and $\mathcal{F}_{p5}$ respectively. The rotational axes corresponding to $q_{j1}$ and $q_{j2}$ are given by the unit vectors, expressed in frame $\mathcal{F}_{b}$, $e_{j1}=(\,\,1\,\,0\,\,0\,\,)^{T}$ and $e_{j2}=R_{x_{b}}(q_{j1})\,(\,\,0\,\,1\,\,0\,\,)^{T}$, respectively, where $R_{k}(*)$ is a pure rotation around axis $k$ by the angle *. Moreover, we assume that each joint composing the passive universal joint $J_{1j}$, with $j=2,3,4$, has viscous friction torques applied about $e_{j1}$ and $e_{j2}$ and symbolized by $\tau_{f_{j1}}$ and $\tau_{f_{j2}}$, respectively.
For the sake of model simplicity, the actuated joint $J_{15}$ is assumed to be driven by two fast servo DC motors taking the joint velocity $u_{v}\in\mathbb{R}^{2}$ as a control input. The motor dynamics is assumed negligible. Furthermore, to improve the formal tractability of the model, we assume that the mass distribution of $B_{5}$ is concentrated at its CoM point which is coincident with its center of rotation. Under such convenient assumptions, the angular moment of inertia of link $B_{5}$ referred to $\mathcal{F}_{p5}$ is zero. This causes the dynamics of the link $B_{5}$ projected onto the minimal coordinates $q_{5}$ to be exactly zero due to the fact that both the time rate of change of the body momentum along with the gravitational wrench vanish when mapped through the screw encoding the kinematical constraints between $B_{1}$ and $B_{5}$. Therefore, the dynamics of $q_{5}$ is solely governed by the motor rotational speed, i.e., $\dot{q}_{5}=u_{v}$
Figure 2: The connection between $B_{j}$ ($j=2,3,4$) and the base $B_{1}$ in PJD2 consists of universal passive joint with a Pulley-Belt-Spring module. $\bar{B}_{j}$ is the intermediate link connecting $B_{1}$ to $B_{j}$. $b$, $p$ and $s$ denote the belt, pulley and a torsion spring, respectively. $\tau_{sj}$ is the torsion spring torque around the rotation axis of $q_{j1}$, i.e. $e_{j1}$.
We differentiate between two types of Passive Joints Designs (PJD), denoted in the following as PJD1 and PJD2, respectively. The design PJD1 has no additional feature with respect to the one described so far, whereas the design PJD2 includes an additional Pulley-Belt-Spring module as illustrated in Fig. 2. This schematic shows a pulley $(p_{j}1)$ transferring the rotation around the axis $\bar{y}_{p_{j}}$, attached to the link ${B}_{j}$, with $j=2,3,4$, by an amount of $q_{j2}$, to equivalent rotation around the axis $\bar{x}_{b}$, attached to the intermediate link $\bar{B}_{j}$, through the belt $(b)$ and pulley $(p_{j}2)$. The torsion spring $(s)$, wound along the axis $\bar{x}_{b}$, is connected to the pulley $(p_{j}2)$ such that it compresses with the rotation of this pulley. The spring is fixed at the other end to a point in $\bar{B}_{j}$.
The spring thus applies a moment to the link $\bar{B}_{j}$ around $\bar{x}_{b}$ that is a function of $q_{j2}$. The entire structure of PJD2 moves with $\bar{B}_{j}$, hence the rotation of joint $q_{j1}$ between ${B}_{1}$ and $\bar{B}_{j}$ has no effect on the operation of the Pulley-Belt-Spring module whatsoever. In other words, the mechanism is unidirectional, meaning that rotation of ${B}_{j}$ around $e_{j2}$ causes a moment applied to $\bar{B}_{j}$ around $e_{j1}$ while the reverse is not true due to the module being fixed in place w.r.t $\bar{B}_{j}$ and housed in it. The rationale behind introducing PJD2 and how the spring is tuned shall become clear in Sec III-A. The analysis presented herein applies to both cases, unless stated otherwise explicitly.
II-B Equations of Motion
We denote with $\mathcal{Q}=(x_{1},q_{5})\in SE(3)\times\mathbb{T}^{8}$ the configuration of the whole system, where $x_{1}\in SE(3)\times\mathbb{T}^{6}$ is defined as
$x_{1}=\left(\begin{array}[]{c}x_{11}\\ x_{12}\end{array}\right):=\left(\begin{array}[]{c}C_{1}^{0}\\ q_{t}\\ \end{array}\right)$ (1)
where $x_{11}:=C_{1}^{0}\in SE(3)$ is the configuration of the base $B_{1}$ relative to the inertial world frame $\mathcal{F}_{w}$ whereas $x_{12}:=q_{t}=(q_{2},q_{3},q_{4})\in\mathbb{T}^{6}$. The velocity of $x_{1}(t)$ is denoted by $x_{2}\in se(3)\times\mathbb{R}^{6}$ which is given as
$x_{2}:=\left(\begin{array}[]{c}x_{21}\\ x_{22}\end{array}\right)=\left(\begin{array}[]{c}V_{1}^{0}\\ \dot{q}_{t}\end{array}\right)$ (2)
where $x_{21}:=V_{1}^{0}=\dot{C}^{0}_{1}(C^{0}_{1})^{-1}\in se(3)$ represents the right differential on $SE(3)$, or equivalently, the right invariant (spatial) twist of the base $B_{1}$ relative to, and expressed in, the inertial world frame $\mathcal{F}_{w}$, while $x_{22}:=\dot{q}_{t}=(\dot{q}_{2},\dot{q}_{3},\dot{q}_{4})\in\mathbb{R}^{6}$ are the joint speeds of the passive joints.
The motion is then described by the dynamical equations
$\begin{split}&M(x_{1})\dot{x}_{2}+h(x_{1},{x}_{2})+\bar{g}(x_{1})-\bar{\tau}_{f}(x_{22})=W^{0}_{app}(x_{1})\\ &\dot{q}_{5}=u_{v}\end{split}$ (3)
where the meaning of all the symbols not yet defined and appearing in (3) are provided in Table I with $j=2,3,4$ and $i=1,2$.
Symbol Definition Formula
$q_{j}$ Joint $J_{1j}$’s configuration $q_{j}=\operatorname{col}(q_{ji})\in\mathbb{T}^{2}$
$q_{t}$ Passive joints configuration $q_{t}=\operatorname{col}(q_{j})\in\mathbb{T}^{6}$
$S_{j}$ Joint $J_{1j}$’s spatial instantaneous screw $S_{j}=(\,S_{j1}\;S_{j2}\,),\ \text{with }S_{ji}\in\mathfrak{se}(3)$
$S_{t}$ Passive joints spatial instantaneous screw list $S_{t}=\operatorname{blockdiag}(S_{j})$
$M$ Generalized mass matrix $M=\begin{pmatrix}M_{bb}&M_{bt}\\ M_{bt}^{T}&M_{tt}\end{pmatrix}$
$M_{bb}$ Generalized base $B_{1}$ mass matrix $M_{bb}=M^{0}_{B_{1}}+\sum^{5}_{j=2}M^{0}_{B_{j}}$, with $M^{0}_{B_{j}}$ being $B_{j}$’s inertia tensor referred to the world frame $\mathcal{F}_{w}$
$M_{bt}$ Generalized coupling mass matrix $M_{bt}=\sum^{4}_{j=2}M^{0}_{B_{j}}S_{j}$
$M_{tt}$ Generalized joint mass matrix $M_{tt}=\operatorname{blockdiag}(S_{j}^{T}M^{0}_{B_{j}}S_{j})$
$\dot{M}^{0}_{j}$ Time derivative of $B_{j}$’s inertia tensor referred to the world frame $\mathcal{F}_{w}$ $\dot{M}^{0}_{j}=-M^{0}_{j}\operatorname{ad}_{V^{0}_{j}}-\operatorname{ad}^{T}_{V^{0}_{j}}M^{0}_{j}$, with $\operatorname{ad}_{V^{0}_{j}}$ denoting the adjoint operator on $V^{0}_{j}$ (the spatial twist of body $B_{j}$), $V^{0}_{j}=V^{0}_{1}+S_{j}\dot{q}_{j}\ \forall j\in\{2,3,4\}$. Also, $\dot{M}^{0}_{5}=-M^{0}_{5}\operatorname{ad}_{V^{0}_{1}}-\operatorname{ad}^{T}_{V^{0}_{1}}M^{0}_{5}$.
$\dot{S}_{j}$ Time derivative of screw $S_{j}$ $\dot{S}_{j}=\operatorname{ad}_{V^{0}_{j}}S_{j}$
$h$ Coriolis & centrifugal terms $h=\big(h_{b}\;\;\operatorname{col}(h_{j})\big)^{T}$
$h_{b}$ Base $B_{1}$ Coriolis & centrifugal terms $h_{b}=\dot{M}_{bb}V^{0}_{1}+\sum^{4}_{j=2}\big(M^{0}_{B_{j}}\dot{S}_{j}+\dot{M}^{0}_{B_{j}}S_{j}\big)\dot{q}_{j}$
$h_{j}$ $B_{j}$’s Coriolis & centrifugal terms projected onto joint $J_{1j}$ $h_{j}=S_{j}^{T}\big((M^{0}_{B_{j}}\dot{S}_{j}+\dot{M}^{0}_{B_{j}}S_{j})\dot{q}_{j}+\dot{M}^{0}_{B_{j}}V^{0}_{1}\big)$
$W_{grav,B_{j}}^{0}$ Spatial gravitational wrench applied at $B_{j}$’s CoM $W_{grav,B_{j}}^{0}=M_{B_{j}}^{0}G^{0}$, with $G^{0}=\operatorname{col}(0_{5},-g)$. Here $g$ is the gravity acceleration.
$\bar{g}$ Gravitational terms $\bar{g}=\big(g_{b}\;\;g_{t}\big)^{T}$
$g_{b}$ Base $B_{1}$ gravitational force $g_{b}=-\sum^{5}_{j=1}W_{grav,B_{j}}^{0}$
$g_{t}$ Joint gravitational torque $g_{t}=-S_{t}^{T}\operatorname{col}(W_{grav,B_{j}}^{0})$
$W^{0}_{app,b}$ Applied external spatial wrench at base $B_{1}$’s CoM $W^{0}_{app,b}=\sum^{5}_{j=2}W_{app,B_{j}}^{0}\in\mathfrak{se}(3)^{*}$, with $W_{app,B_{j}}^{0}\in\mathfrak{se}(3)^{*}$ being the applied spatial wrench at $B_{j}$’s CoM
$W^{0}_{app,t}$ Applied external spatial wrench at pendulum links $B_{2},B_{3},B_{4}$ projected onto the respective joints through $S_{t}$ $W^{0}_{app,t}=S_{t}^{T}\operatorname{col}(W_{app,B_{j}}^{0})$
$W^{0}_{app}$ Total applied external wrenches $W^{0}_{app}=\big(W^{0}_{app,b}\;\;W^{0}_{app,t}\big)^{T}$
$\tau_{fj}$ Joint $J_{1j}$’s applied viscous friction torque $\tau_{fj}=\operatorname{col}(\tau_{f_{ji}})\in\mathbb{R}^{2}$, and $\bar{\tau}_{f}=\big(0_{6\times 1}\;\;\operatorname{col}(\tau_{f_{j}})\big)^{T}$
TABLE I: Variable definitions and formulas
Denote the full state vector $x(t)$ as $x=(x_{1},x_{2},{q}_{5})^{T}\in\mathbb{X}\equiv SE(3)\times\mathbb{T}^{8}\times se(3)\times\mathbb{R}^{6}$. Then (3) can be put in the state space form $\Sigma_{1}$ as follows
[…]
where $[.]$ denotes the matrix form of $se(3)$, and $D_{b}$, $D_{t}$ and the input vector $u$ are defined as
$\begin{split}D_{b}=&\left(\begin{array}[]{cccc}\operatorname{Ad}^{T}_{(C^{0}_{f_{2}})^{-1}}\bar{f}_{2}&\operatorname{Ad}^{T}_{(C^{0}_{f_{3}})^{-1}}\bar{f}_{3}&\operatorname{Ad}^{T}_{(C^{0}_{f_{4}})^{-1}}\bar{f}_{4}&\operatorname{Ad}^{T}_{(C^{0}_{f_{5}})^{-1}}\bar{f}_{5}\end{array}\right)_{6\times 4}\\ D_{t}=&\left(\begin{array}[]{ccc}S_{2}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{2}})^{-1}}\bar{f}_{2}&0&0\\ 0&S_{3}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{3}})^{-1}}\bar{f}_{3}&0\\ 0&0&S_{4}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{4}})^{-1}}\bar{f}_{4}\end{array}\right)_{6\times 3}\\ u=&\left(\begin{array}[]{cccccc}f_{2}&f_{3}&f_{4}&f_{5}&u_{v_{1}}&u_{v_{2}}\end{array}\right)_{6\times 1}^{T}\in\mathbb{R}^{6}\end{split}$
[…]
Result 1 (Input dimension property).
The proposed design satisfies the input dimension property. In fact the dimension of the manifold of the configuration of the base frame $C_{1}^{0}$ is six and the dimension of the input is $\operatorname{dim}(u)=6$ (four propeller speeds and two servomotor speeds). In other words, the system is square with respect to the output being $C_{1}^{0}$, the pose of $B_{1}$.
III Study of Attainable Equilibria and
Internal Forces at each Equilibrium
In this section we obtain the set of forced equilibria of the dynamics (4) and investigate their properties in terms of the corresponding four thrust forces produced at these configurations. We begin by giving some preliminary definitions. Afterwards, the main result is summarized in Proposition 1 whose proof is found in Sec. VII-B.
The value of a variable at an equilibrium is identified by a subscript $e$. Imposing $\dot{x}{(t)}=0$ in (4) implies $({x}_{2}(t),\dot{q}_{5}(t))=0$, which defines an equilibrium condition at which the state and the input become, respectively, $x_{e}=(x_{1e},0,{q}_{5e})^{T}$ and $u_{e}=(f_{e}\,\,u_{ve})^{T}$ with $f_{e}=(f_{2e}\,\,f_{3e}\,\,f_{4e}\,\,f_{5e})^{T}$ and $u_{ve}=(u_{v_{1e}}\,\,u_{v_{2e}})^{T}$. This condition yields $h(x_{1}(t),{x}_{2}(t))=0$ since it depends quadratically on the configuration velocity state ${x}_{2}(t)$, as apparent from Table I. Similarly, the viscous friction $\bar{\tau}_{f}(x_{22}(t))=0$. For PJD1, this reduces (4) to this static balance vector equation whose admissible solutions are the equilibrium state and input pairs $(x_{e},u_{e})$ such that
$g_{b}(C^{0}_{1e},q_{te})=W^{0}_{app,b}(f_{e},q_{te},q_{5e}),$ (6)
$g_{t}(C^{0}_{1e},q_{te})=W^{0}_{app,t}(q_{te},f_{2e},f_{3e},f_{4e}):=(\,\,D_{t}\,\,0_{6\times 3}\,\,)u_{e}$
$\quad\quad\quad\quad\quad=(\,\,\tau_{d2}\,\,0\,\,\tau_{d3}\,\,0\,\,\tau_{d4}\,\,0\,\,)_{6\times 1}^{T},\quad 0_{2}=u_{ve}$ (7)
where $\tau_{dj}=k_{j}f_{je}\sin(q_{j2e})$ – with $j=2,3,4$ – represents the drag component of the propeller of $B_{j}$ factorized on the axis $e_{j1}$ of the joint variable $q_{j1}$ at an equilibrium; $C^{0}_{1e}$ and $q_{te}=\operatorname{col}(\,q_{je})$, with $q_{je}=\operatorname{col}(\,q_{jie})$ $\forall j=2,3,4\text{ and }\forall i=1,2$, symbolizing an equilibrium of the base configuration $C^{0}_{1}$ and the passive joints configuration $q_{t}$, respectively.
III-A Effect of the Aerodynamic Drag Torque at the Equilibrium
The presence of the torque $\tau_{dj}$ in (7) leads to the important consequence that the value of the joints $q_{te}$ at the equilibrium is not solely determined by the gravitational torque at equilibrium $g_{te}$, as it is the case for the 2D MAV presented in  [6], but also by a component of the equilibrium thrust $f_{e}$ at equilibrium.
This residual torque can be eliminated at any equilibrium by the spring introduced in PJD2. In fact, we can tune each torsion spring to have a variable stiffness $k_{s}:=\frac{\partial\tau_{sj}}{\partial q_{j2}}=k_{j}f_{je}\cos(q_{j2})$ where $\tau_{sj}$ is the torque absorbed by the spring at the joint angle $q_{j2}$. This task is facilitated by the fact that the term $k_{j}f_{je}$ is always constant since $f_{je}$ is found to be constant for any equilibrium of PJD2, as proven in Sec. VII-B (49). As a result, this component $\tau_{dj}$ is canceled at any PJD2 equilibrium, i.e., $\tau_{dj}=-\tau_{sje}$, transforming (7) to
$g_{t}(C^{0}_{1e},q_{te})=0_{6},\,\,\,u_{ve}=0_{2}.$ (8)
We solve (7) and (8) in Sec. VII-B for the joint equilibrium $q_{je}$. To distinguish between a joint equilibrium of PJD1 and PJD2, let $q_{te}^{[i]}$ denote a joint equilibrium solution of PJDi, with $i=1,2$, and $f^{[i]}_{e}$ be the associated thrust, respectively.
III-B Minimal Thrust at Equilibrium vs Presence of Internal Forces
Before presenting the result on the equilibria property of the proposed MAV (4) we provide two definition for the thrust magnitudes at the equilibrium, in the case in which they are minimal and in the case in which there is internal force.
Definition 1.
A vector of thrust magnitudes $f\in\mathbb{R}_{\geq 0}^{l}$, where $\mathbb{R}_{\geq 0}^{l}:=\left\{s=col(s_{i})\in\mathbb{R}^{l}\mid s_{i}\geq 0,\ \forall i\in\{1,\dots,l\}\right\}$, is said to be minimal at the equilibrium if it belongs to $\mathcal{F}_{\rm min}^{\rm eq}$, defined as:
$\mathcal{F}_{\rm min}^{\rm eq}:=\bigg\{f=col(f_{i})\in\mathbb{R}_{\geq 0}^{l}\bigg|\,\sum^{l}_{i=1}{f}_{i}=m_{\rm tot}\,g\bigg\}$ (9)
where $m_{\rm tot}$ is the total mass of the system. Instead, $f$ is said to have internal forces at the equilibrium if it belongs to $\mathcal{F}_{\rm int.for.}^{\rm eq}$, defined as:
$\mathcal{F}_{\rm int.for.}^{\rm eq}:=\bigg\{f=col(f_{i})\in\mathbb{R}_{\geq 0}^{l}\bigg|\,\sum^{l}_{i=1}{f}_{i}>m_{\rm tot}\,g\bigg\}$ (10)
A MAV design whose thrust magnitudes are minimal at all its equilibria is said to have no internal forces at its equilibria.
In other words, a minimal thrust magnitudes vector $f\in\mathcal{F}_{\rm min}^{\rm eq}$ has the least amount of total thrust magnitudes necessary to counteract the gravity effect at the equilibrium, whereas $f\in\mathcal{F}_{\rm int.for.}^{\rm eq}$ means that the vehicle generates more total thrust than what is strictly necessary for compensating gravity, thus having some wasted internal force.
Moreover, there exists a unique orientation for each propeller in the system such that that when all the propellers assume such orientation the absence of internal force is guaranteed at hovering. The following lemma provides the formal statement.
Lemma 1.
Consider a vehicle with a thrust magnitudes vector $f=col(f_{i})$ $\in\mathbb{R}_{\geq 0}^{l}$ with the corresponding propellers configurations $C^{0}_{i}(C^{0}_{1},\alpha_{i})\in SE(3)$, relative to a right-handed inertial frame whose z-axis direction opposes the axis along which the gravitational force is applied, where $C^{0}_{1}\in SE(3)$ is the base configuration and $\alpha_{i}\in\mathbb{S}^{2}$ is the unit direction vector of the propeller $i$ expressed in the inertial frame. At any equilibrium for the MAV with a given $\bar{C}^{0}_{1}$, there is no internal force generated at such equilibrium if and only if all propellers, whose thrust magnitudes are $f_{i}>0$, are pointing in the direction of inertial frame’s z-axis, i.e. $C^{0}_{i}(\bar{C}^{0}_{1},\alpha_{i})\,e_{3}=e_{3}$ $\forall i\in\{1,\dots,l\}$ with $e_{3}=(\,0\,0\,1\,0\,)^{T}$.
Proof.
See Sec. VII-A. ∎
The following proposition establishes the main result of this section.
Proposition 1.
Consider any pose for the base $B_{1}$ denoted by $C^{0}_{1e}$. Then the following statements hold true and are equivalent for the system (4):
  1. 1.
    For any base pose $C^{0}_{1}\in SE(3)$, there exists an equilibrium for PJD2 in which the joint equilibrium configuration $q_{te}^{[2]}$ satisfies $C^{0}_{j}(C^{0}_{1e},q^{[2]}_{je})e_{3}=e_{3}$ $\forall j\in\{2,\dots,5\}$ and a thrust magnitudes vector $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ is produced whose elements are constant $\forall C^{0}_{1}$ and equal to (49).
  2. 2.
    For any base pose $C^{0}_{1}$, PJD2 has no internal force at any equilibrium $q_{te}^{[2]}$, by Definition 1, as a corollary of $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ (Statement 1).
  3. 3.
    When $\tau_{dj}\neq 0$, PJD1 has no joint equilibrium in which each propeller produces a minimal thrust: $C^{0}_{j}(C^{0}_{1e},q^{[1]}_{je})e_{3}\neq e_{3},\,\forall\tau_{dj}\neq 0$.
  4. 4.
    When $\tau_{dj}\neq 0$, it follows from Statement 3 and Lemma 1 that PJD1 has an internal force $f^{[1]}_{e}\in\mathcal{F}_{\rm int.for.}^{\rm eq}$. The amount of internal force changes with the parameters $k_{i}$,c\,c and $m_{p}$, defined in Table II.
Proof.
See Sec. VII-B. ∎
The state equilibria sets for PJD1 and PJD2 can thus be defined as $\mathbb{D}^{[1]}_{x_{e}}$ and $\mathbb{D}^{[2]}_{x_{e}}$, respectively, with $i=1,2$
[…]
while the associated thrust equilibrium sets are given by
[…]
We conclude the section remarking that the perfect vertical alignment of propellers in any PJD2 equilibrium is possible thanks to each spring torque $\tau_{sje}$ absorbing the residual drag $\tau_{dj}$ affecting the joint variable $q_{j1}$. This allows any joint equilibrium $q_{te}^{[2]}$ to be dictated only by the gravity acting on the CoMs of the pendulum-like links B2-B4, given that the pivots, i.e., the joints, are passive.
Remark 1.
The design using PJD1’s constitutes still an excellent compromise between mechanical simplicity and energy efficiency, with the ability to outperform the state of the art designs when it comes to the amount of internal force needed at the the equilibria. In fact, for typical values of the aerodynamic and kinematic parameters the difference in the amount of internal forces generated in PJD1 design remains small compared to PJD2 case, as it can be seen for example in the simulations of Section V. This supports the viability of the concept from a mechanical perspective by using the simpler design PJD1, if one does not want to resort to the particular joint design in PJD2.
IV Stabilization of the Equilibria
This section presents the solution of the problem of whether every equilibrium point in the set $\mathbb{D}^{[i]}_{x_{e}}$, given in (11), can be made closed-loop asymptotically stable, at least locally, by an action of a regular state feedback law. For clarity, we first provide an informal outline of the solution’s derivation and establish necessary definitions, leading to the core findings stated in Theorem 1 and Proposition 2.
IV-A Informal Outline of the Solution
Denote with $\Sigma_{i}(x;u)$ a system $\Sigma_{i}$ whose states and inputs are respectively $x$ and $u$. The analysis departs by noticing that there exists no static state feedback control $u=\alpha(x)$ rendering the closed-loop system $\Sigma_{1}(x;\alpha(x))$ in (4) locally asymptotically stable (LAS) in an open neighborhood $\mathbb{U}_{{x}_{e}}\subseteq\mathbb{X}$ of an equilibrium ${x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}$, since it turns out that $\Sigma_{1}$ does not possess a well-defined vector relative degree at any $x_{e}\in\mathbb{D}^{[i]}_{x_{e}}$  [22]. However, the goal of stabilizing this closed-loop system at $x_{e}$ is still attainable by means of dynamic state feedback.
Towards this end, we first observe that $\Sigma_{1}$ in (4) can be decomposed into two subsystems
$\Sigma_{1}:\Sigma_{2}(x_{12},x_{2},q_{5};u,x_{11})\text{ and }\Sigma_{3}(x_{11};x_{21})$ (13)
where $\Sigma_{2}$ is found to belong to the class of dynamically input/output feedback linearizable systems w.r.t an output taken from $V_{1}^{0}=:x_{21}$. Second, by applying the dynamic extension algorithm [23] on $\Sigma_{2}$, an extended system $\Sigma_{\bar{2}}(\bar{x};w,x_{11})$ results, where the extended state is $\bar{x}\in\bar{\mathbb{X}}\equiv\mathbb{T}^{8}\times se(3)\times\mathbb{R}^{10}$ and $w\in\mathbb{R}^{6}$ is the new input for its input precompensator. For this extended system $\Sigma_{\bar{2}}$, we compute a local diffeomorphism $H(\bar{x}):\mathbb{U}_{\bar{x}_{e}}\rightarrow\mathbb{U}_{\tilde{x}_{e}}$, with a domain in $\mathbb{U}_{\bar{x}_{e}}\subseteq\bar{\mathbb{X}}$ of $\bar{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}\cup\mathbb{D}^{[i]}_{f_{e}}$ (12), transforming it into its normal form represented by $\Sigma^{H}_{\tilde{2}}(\zeta,\eta;w,x_{11})$.
This normal form $\Sigma^{H}_{\tilde{2}}(\tilde{x};w,x_{11})$ has a state $\tilde{x}:=(\zeta,\eta)\in\mathbb{U}_{\tilde{x}_{e}}\subseteq\tilde{\mathbb{X}}\equiv se(3)\times se(3)\times\mathbb{T}^{6}\times\mathbb{R}^{6}$, where $\mathbb{U}_{\tilde{x}_{e}}$ denotes an open neighborhood of $\tilde{x}_{e}=:({\zeta_{e}},{\eta_{e}})=H(\bar{x}_{e})$. The normal form is a feedback interconnection between the dynamics of the I/O map $\Sigma_{\zeta}(\zeta;\eta,w,x_{11})$ and some internal dynamics $\Sigma_{\eta}(\eta;\zeta,x_{11})$. $\Sigma_{\zeta}$ can be feedback linearized, becoming $\Sigma_{\zeta}^{w}(\zeta;v)$, by an action of a regular static state feedback $w=\alpha_{w}(\tilde{x},x_{11},v)$ with a domain in $\mathbb{U}_{\tilde{x}_{e}}\times SE(3)\times se(3)$, where $v\in se(3)$ is a virtual input.
At this point, we return to the $\Sigma_{1}$ decomposition in (13). Let us denote with $\Sigma^{H}_{3}(x_{11};\zeta)$ the system $\Sigma_{3}$ in (13) after the diffeomorphism $H$ is applied and with $\Sigma_{\bar{\zeta}}(\bar{\zeta};v)$ its composition with the feedback linearized system $\Sigma_{\zeta}^{w}$, i.e. $\Sigma_{\bar{\zeta}}$ is given by
$\Sigma_{\bar{\zeta}}:\Sigma_{\zeta}^{w}(\zeta;v)\text{ and }\Sigma^{H}_{3}(x_{11};\zeta)$ (14)
where $\bar{\zeta}$ denotes $\zeta$ extended with $x_{11}$. We can now conclude that if there exists a static state feedback $v=\alpha_{v}(\bar{\zeta})$ locally asymptotically stabilizing $\Sigma_{\bar{\zeta}}$ at $\bar{\zeta}_{e}:=(\zeta_{e},x_{11e})$, then, from Theorem 10.3.1 in [22], the cascaded interconnection of the autonomous system $\Sigma_{\bar{\zeta}}(\bar{\zeta};\alpha_{v}(\bar{\zeta}))$ and $\Sigma_{\eta}(\eta;\bar{\zeta})$ is guaranteed to be LAS in $\mathbb{U}_{\tilde{x}_{e}}\times SE(3)$ of $(\tilde{x}_{e},x_{11e})$, provided that the zero dynamics $\Sigma_{\eta}(\eta;0)$ is LAS at $\eta_{e}$}.
This means that, as long as trajectories of $\Sigma_{\bar{\zeta}}$ and $\Sigma_{\eta}$ remain in $\mathbb{U}_{\tilde{x}_{e}}\times SE(3)$, any base pose $x_{11e}$ of the original system $\Sigma_{1}$ (4) can be made locally asymptotically stable by a dynamic state feedback.
IV-B Main Results
Theorem 1 formalizes this result while Proposition 2 provides the stabilizing feedback $v=\alpha_{v}(\bar{\zeta})$ whose construction details, drawing upon a backstepping strategy on $\Sigma_{\bar{\zeta}}$, are laid out in the proof in Sec. VII-D.
Theorem 1.
Let the original system $\Sigma_{1}$ (4) be decomposed as in (13) and define the system $\Sigma_{3}(x_{11};x_{21})$ to be $\Sigma_{3}:\dot{x}_{11}=[{x}_{21}]\,{x}_{11}$ while $\Sigma_{2}(x_{12},x_{2},q_{5};u,x_{11})$ is
$\Sigma_{2}:\left\{\begin{aligned} \dot{x}=&\left(\begin{array}[]{c}x_{22}\\ M^{-1}(-h-g+\bar{\tau}_{f})\\ 0_{2\times 1}\end{array}\right)+\left(\begin{array}[]{c}0_{6\times 6}\\ M^{-1}\left(\begin{array}[]{cc}D_{b}&0_{6\times 2}\\ D_{t}&0_{6\times 3}\\ \end{array}\right)\\ \begin{array}[]{cc}0_{2\times 4}&I_{2\times 2}\end{array}\end{array}\right)u\\ =&F(x)+G(x)u\end{aligned}\right$.
Then, these statements are true
  1. 1.
    $\Sigma_{2}$ is dynamically input/output feedback linearizable w.r.t the I/O pair $(u,h_{y})$, in an open neighborhood $\mathbb{U}_{\bar{x}_{e}}\subseteq\bar{\mathbb{X}}$ of $\bar{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}\cup\mathbb{D}^{[i]}_{f_{e}}$, where the output function $h_{y}(t):t\rightarrow se(3)$ is $h_{y}(t)=V_{1}^{0}=:x_{21}$. This is accomplished by the input precompensator
    $z(t)=col(z_{i}):=f(t)\in\mathbb{R}^{4},\dot{z}=\bar{w}\in\mathbb{R}^{4},i=2,\ldots,5$ (16)
    and the linearizing state feedback ${w}:=(\bar{w},u_{v_{1}},u_{v_{2}})^{T}$,
    ${w}=D_{2}^{-1}M_{b}(-F_{2}+v)$ (17)
    where $v\in se(3)$ is the virtual input, and $F_{2}$, $D_{2}$ and $M_{b}$ are given in (68), (61) and (66). This path-connected open neighborhood $\mathbb{U}_{\bar{x}_{e}}$ of $\bar{x}_{e}$ is defined as
    $\mathbb{U}_{\bar{x}_{e}}=\Big\{\bar{x}(t)\in\bar{\mathbb{X}}\,\Bigr\rvert\,\operatorname{det}\big(D_{2}(\bar{x})\big)\neq 0\Big\}$ (18)
  2. 2.
    For any base pose $x_{11}\in SE(3)$, The original system $\Sigma_{1}$, extended with the precompensator (16), has locally, in $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$, asymptotically stable (LAS) zero dynamics w.r.t the output $h_{y}$, iff $\forall\bar{x}\in\mathbb{Z}_{x}-\{\bar{x}_{e}\}$
    $\|\tilde{f}\|<\frac{1}{\|D_{3}(q_{t}-q_{te})\|}\lambda_{\text{min}}(D_{f})\|\dot{q}_{t}\|$ (19)
    where $\tilde{f}=col({f}_{i})\,,i=2,3,4$, $\mathbb{Z}_{x}$ is a set defined in (77), $D_{3}$ is given in (71) and $D_{f}>0\in\mathbb{R}^{6\times 6}$ contains the constant passive joints friction coefficients.
Proof.
See Sec. VII-C
Proposition 2.
Consider the virtual control input $v$ given by
$v=\ddot{\zeta}_{1e}+\bar{K}_{d}(\dot{\zeta}_{1e}-\dot{V}^{0}_{1})+\bar{K}_{a}({\zeta}_{1e}-V^{0}_{1})$ (20)
where $\zeta_{1e}$, $\dot{\zeta}_{1e}$, $\ddot{\zeta}_{1e}$ $\in se(3)$ are respectively given in (92) and (104). $\bar{K}_{d}$ and $\bar{K}_{a}$ $\in\mathbb{R}^{6\times 6}$ are positive definite diagonal gain matrices chosen such that to make (ABK)(A-BK) Hurwitz where
$A=\left(\begin{array}[]{c}0_{6\times 6}\quad I_{6\times 6}\\ 0_{6\times 6}\quad 0_{6\times 6}\\ \end{array}\right),\,B=\left(\begin{array}[]{c}0_{6\times 6}\\ I_{6\times 6}\\ \end{array}\right),\,K=\left(\begin{array}[]{c}\bar{K}_{a}\quad\bar{K}_{d}\end{array}\right)$
Given the precompensator in (16) and the linearizing control law in (17), substituting $v$ into (17) asymptotically stabilizes the extended system $\Sigma_{1}$ with (16) in the open neighborhood $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$, where $\mathbb{U}_{\bar{x}_{e}}\subseteq\bar{\mathbb{X}}$ of $\bar{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}\cup\mathbb{D}^{[i]}_{f_{e}}$ is given in (18).
Proof.
See Sec. VII-D
IV-C Efficient Implementation of the Controller
We introduce here a novel version of the Articulated-body Inertia Algorithm (ABA), originally derived in [24], which is modified to recursively obtain the acceleration $(\ddot{q}_{t},\dot{V}^{0}_{1})$ needed to implement the feedback linearizing law (17) (20) efficiently, instead of inverting the matrix $M$ in (3).
For the system in Fig. 1, a modified ABA computes $(\ddot{q}_{t},\dot{V}^{0}_{1})$ executing the following computations for each $j=2,3,4$
$\begin{split}{M}_{j2}^{A}&={M}^{0}_{B_{j}}\\ \psi_{j2}&=\left(S_{j2}^{T}M^{A}_{j2}S_{j2}\right)^{-1}\\ {W}_{j2}^{A}&=-\operatorname{ad}^{T}_{{V}_{j2}^{0}}{M}^{0}_{B_{j}}{V}^{0}_{j2}-W_{app,B_{j}}^{0}-W_{grav,B_{j}}^{0}\\ {M}_{j1}^{A}&=M_{j2}^{A}-M_{j2}^{A}S_{j2}\psi_{j2}S_{j2}^{T}M_{j2}^{A}\\ \psi_{j1}&=\left(S_{j1}^{T}M^{A}_{j1}S_{j1}\right)^{-1}\\ {W}_{j1}^{A}&=M_{j2}^{A}\left(S_{j2}\ddot{\tilde{q}}_{j2}+\dot{S}_{j2}\dot{q}_{j2}\right)+{W}_{j2}^{A}\end{split}$ (22)
where $\ddot{\tilde{q}}_{ji}$, with $i=1,2$, is given by:
$\begin{split}\ddot{\tilde{q}}_{ji}&=\psi_{ji}\left(\tau_{f_{ji}}-S_{ji}^{T}\left(M^{A}_{ji}\dot{S}_{ji}\dot{q}_{ji}+{W}_{ji}^{A}\right)\right)\end{split}$ (23)
Afterwards, the articulated inertia ${M}_{1}^{A}$ and bias wrench ${W}_{1}^{A}$ for the base are computed as
$\begin{split}{M}_{1}^{A}&={M}^{0}_{B_{1}}+{M}^{0}_{B_{5}}+\sum_{j=2}^{4}\left(M_{j1}^{A}-M_{j1}^{A}S_{j1}\psi_{j1}S_{j1}^{T}M_{j1}^{A}\right),\\ {W}_{1}^{A}&=-\operatorname{ad}^{T}_{{V}_{1}^{0}}({M}^{0}_{B_{1}}+{M}^{0}_{B_{5}}){V}^{0}_{1}-W_{app,B_{1}}^{0}-W_{grav,B_{1}}^{0}-W_{grav,B_{5}}^{0}\\ &\quad+\sum_{j=2}^{4}\big(M_{j1}^{A}(S_{j1}\ddot{\tilde{q}}_{j1}+\dot{S}_{j1}\dot{q}_{j1})+{W}_{j1}^{A}\big).\end{split}$
Hence, the following six linear equations are solved for the base acceleration twist $\dot{V}_{1}^{0}$ using $LDL^{T}$ method, since $M_{1}^{A}$ is symmetric positive definite
$0=M_{1}^{A}\dot{V}_{1}^{0}+{W}_{1}^{A}.$ (25)
The obtained value of $\dot{V}_{1}^{0}$ is then used to retrieve the joint accelerations by evaluating the following expressions:
$\begin{split}\ddot{q}_{j1}&=-\psi_{j1}S_{j1}^{T}M^{A}_{j1}\dot{V}_{1}^{0}+\ddot{\tilde{q}}_{j1}\\ \dot{V}_{j1}^{0}&=\dot{V}_{1}^{0}+S_{j1}\ddot{q}_{j1}+\dot{S}_{j1}\dot{q}_{j1}\\ \ddot{q}_{j2}&=-\psi_{j2}S_{j2}^{T}M^{A}_{j2}\dot{V}_{j1}^{0}+\ddot{\tilde{q}}_{j2}.\end{split}$ (26)
Parameter Value Unit Description
$m_{p}$ $0.5$ kg Link $B_{j}$ mass, $j=2,3,4$
$m_{a}$ $0.5$ kg Link $B_{5}$ mass
$m_{b}$ $1.0$ kg Base $B_{1}$ mass
$a$ $0.5$ m Length between any two adjacent joints
$c$ $0.25$ m Length from a joint to the respective $B_{j}$’s CoM
$g$ $9.81$ m/s2 Gravitational acceleration
$I_{px}$ $2.1875\times 10^{-3}$ kg·m2 Link $B_{j}$ inertia about $x_{p}$-axis
$I_{py}$ $2.1875\times 10^{-3}$ kg·m2 Link $B_{j}$ inertia about $y_{p}$-axis
$I_{pz}$ $1.168\times 10^{-6}$ kg·m2 Link $B_{j}$ inertia about $z_{p}$-axis
$I_{bx}$ $8.3333\times 10^{-3}$ kg·m2 Inertia of base $B_{1}$ about $x_{b}$-axis
$I_{by}$ $8.3333\times 10^{-3}$ kg·m2 Inertia of base $B_{1}$ about $y_{b}$-axis
$I_{bz}$ $1.6667\times 10^{-2}$ kg·m2 Inertia of base $B_{1}$ about $z_{b}$-axis
$b_{fx}$ $0.9$ N.m.s/rad Viscous friction coefficient about $e_{j1}$-axis
$b_{fy}$ $0.9$ N.m.s/rad Viscous friction coefficient about $e_{j2}$-axis
$|k|$ $0.01$ m Ratio of drag-trust coeff. for $B_{j}$’s propeller
TABLE II: Plant Parameters with Definitions and Units
V Simulation Tests
This section presents simulation results designed to establish confidence in the generalizability of our findings beyond the standard model used for formal proofs. To achieve this, we evaluate the proposed controller using a more complex model that incorporates non-idealities such as noise and uncertain parameters. To show that the main body $B_{1}$ of the proposed MAV (4), with the state feedback controller in (17) (20) whose dynamics are given in (16), is capable of stabilizing its attitude independently from its position and vice versa in both designs PJD1 and PJD2, two indicative values for the base pose at equilibrium are selected. The attitude set point is chosen such that the MAV stabilizes a base configuration making the desired frame $\mathcal{F}_{b}$ rotated around the inertial frame $y_{w}$-axis by $60^{\circ}$, i.e., $C^{0}_{1e}=\left(\begin{smallmatrix}R^{0}_{1e}&0_{3\times 1}\\ 0_{1\times 3}&1\end{smallmatrix}\right)$ with $R^{0}_{1e}=R_{x}(0)R_{y}(60^{\circ})R_{z}(0)$, whereas the position step reference is formed by the desired $C_{1e}^{0}$ pose having a constant orientation and a position of $\mathcal{F}_{b}$’s origin equal to 11 mm along the $y_{w}$-axis, i.e., $C^{0}_{1e}=\left(\begin{smallmatrix}I_{3\times 3}&(\,0\,1\,0\,)^{T}\\ 0_{1\times 3}&1\end{smallmatrix}\right)$. The position reference is applied for the first 60 seconds of the simulation. Upon convergence to the desired position, the attitude maneuver starts. The corresponding state and input evolutions over time are reported in Fig. 3.
The orientation and position errors of frame $\mathcal{F}_{b}$ are denoted by $e_{\phi_{1,l}}$ and $e_{p_{1,l}}$, respectively, with $l=\{x,y,z\}$. $e_{\phi_{1,l}}$ is defined as $e_{\phi_{1,l}}:=\phi_{e,l}-\phi_{l}$ where $\phi_{e,l}$ and $\phi_{l}$ are the desired and actual angles resulting from transforming the rotation matrices in $C^{0}_{1e}$ and $C^{0}_{1}$ to the Euler angles sequence $XYZ$ during the simulation time, respectively. $e_{p_{1,l}}$ is given by $e_{p_{1,l}}:=p_{e,l}-p_{l}$ where $p_{e}$ and $p$ are the desired and actual position vectors of $\mathcal{F}_{b}$’s origin in the inertial frame, respectively. The $e_{w^{0}_{1}}$ and $e_{v^{0}_{1}}$ are the evolution of the angular and linear velocities components of the base twist $V^{0}_{1}$, respectively.
To better emulate realistic conditions, the simulation in Fig. 3 includes parametric uncertainty and input noise. The noise is a low-pass filtered Gaussian noise with a state ${\beta}(t)\in\mathbb{R}^{6}$ and dynamics
$\dot{{\beta}}(t)=-k{\beta}(t)+{\mu},$ ${\mu}\in\mathcal{N}({0},l^{2}I_{6\times 6}),$ (27)
where $l=0.1$ and $k=45$. Fig. 6 shows the components of the noise $\beta(t)$ added to the actuators during the simulation in Fig. 3. The true parameter values, listed in Table II, are each perturbed by $5\%$. Additionally, the initial conditions are not an admissible equilibrium and given by $(q_{t}(0),q_{5}(0))=0.1\times\boldsymbol{1}_{8\times 1}\,\text{rad}$, ${x}_{2}(0)=0_{12\times 1}$, and $f(0)=1.05\,f^{[2]}_{e}$, with $f^{[2]}_{e}$ calculated from (49). The initial base configuration for simulation in Fig. 3 is $C^{0}_{1}(0)=\left(\begin{smallmatrix}R^{0}_{1}(0)&(\,1\,-1\,-1\,)^{T}\\ 0_{1\times 3}&1\end{smallmatrix}\right)$ with $R^{0}_{1}(0)=R_{x}(-60^{\circ})R_{y}(-30^{\circ})R_{z}(60^{\circ})$, introducing initial attitude and position errors.
(a) PJD1
(b) PJD2
Figure 3: Time evolutions of the state and input for a common simulation are presented for two designs: (a) PJD1 and (b) PJD2. The simulation comprises two phases. First (0-60s), a step reference of 1m along the inertial $y_{w}$ axis is commanded for the base position, with the initial attitude held constant. Second (60-130s, gray-shaded), a step reference demands a $60^{\circ}$ rotation about the inertial $y_{w}$ axis for the attitude, while the position at $t=60$s is maintained. Imperfections such as large initial errors, uncertainty, and noise contribute to the angular velocity and attitude errors during both the initial position maneuver and the subsequent attitude maneuver
Figure 4: Time evolution of the zero dynamics states showing convergence.
Figure 5: Comparison of the total thrust evolution between PJD1 and PJD2 for the same scenario in Fig. 3. In steady state, PJD2 thrust magnitudes sum up to exactly the minimal value (9), unlike its PJD1 counterpart, which is about 2% higher (still much better compared to the internal forces required by state of the art designs).
Figure 6: Evolution of noise states added to the inputs in the simulation in Fig. 3.
To cope with parametric uncertainty, the version of the virtual control including an integral term (108) in Remark 3 is utilized to eliminate the steady state error. The controller gains $\bar{K}_{d}=\operatorname{diag}(5.5576)$ and $\bar{K}_{a}=\operatorname{diag}(14.9433)$ are obtained from applying LQR on the error linear system (107), with weighting matrices on its state $Q=\operatorname{diag}(0.0001)\in\mathbb{R}^{12\times 12}$ and input $R=\operatorname{diag}(0.0001)\in\mathbb{R}^{6\times 6}$, while $K_{I}=\operatorname{diag}(60)$ is selected by trial-and-error to zero the steady-state error. The gain matrix $K_{p}=\operatorname{diag}(0.1)$ appearing in the kinematics controller (92) affects how fast $C^{0}_{1}(t)$ is desired to converge to $C^{0}_{1e}$, which in turn reflects on how fast the reference trajectory $\zeta_{1e}(t)$ in the error system (107) evolves. When the closed-loop error system (107) operates with fixed values of $\bar{K}_{d}$ and $\bar{K}_{a}$, the higher this $K_{p}$ gain is, the larger the error state values become, requiring greater thrusts and faster tilting speed of the actively-tilting propeller.
Furthermore, to show that the zero dynamics (78) in the simulations has a stable behavior as formally proven in Sec. VII-C for the standard model, we simulated its evolution for initial conditions $C^{0}_{1}(0)=I_{4\times 4},\,q_{t}(0)=-30^{\circ}\times\boldsymbol{1}_{6\times 1}\,$, $\dot{q_{t}}(0)=0.1\text{ rad/s }\times\boldsymbol{1}_{6\times 1}$ in the nominal settings. The results in Fig. 4 indicate convergence to zero of the passive joint variables and corresponding speeds, which is indeed the corresponding joint equilibrium in $\bar{x}_{e}$ for the base pose $C_{1e}^{0}=C^{0}_{1}(0)=I_{4\times 4}$.
We conclude by highlighting the main difference between PJD1 and PJD2 by comparing the total thrust magnitudes for the scenario in Fig. 3. It can be observed from Fig. 5 that PJD1 equilibrium thrust satisfies $\sum^{5}_{j=2}f^{[1]}_{je}>\sum^{5}_{j=2}f^{[2]}_{je}$ with $f^{[2]}_{e}=f_{min}\in\mathcal{F}_{\rm min}^{\rm eq}$. In fact, for this desired $C_{1e}^{0}$, the residual torque is $\tau_{dj}\neq 0$ and thus the corresponding joint equilibrium $q_{t}^{[1]}$ yields an equilibrium thrust $f_{e}^{[1]}$ satisfying $\sum^{5}_{j=2}f^{[1]}_{je}\approx 1.02\,f_{min}$, hence witnessing about $2\%$ increase over the minimal value $f_{e}^{[2]}$, which is achieved in PJD2, reinforcing the findings in Proposition 1.
Videos for the numerical simulations in Fig. 3 are available at https://tinyurl.com/bdur26au.
VI Conclusion
In this work, we have presented a novel MAV design idea that attains omnidirectionality in the 3D space with no internal force at the equilibrium and with minimum number of actuators. This is accomplished by first introducing the proposed vehicle’s dynamical model composed of a main body (base), whose configuration evolves on $SE(3)$ thus avoiding the singularity issues stemming from using local coordinates for the configuration manifold, and 4 propeller-carrying links connected to this base by 4 universal joints, 3 of which are passive.
Afterward, the equilibrium set is derived. This vehicle has an equilibrium for the joint positions at every pose of the base. If the universal joints are additionally equipped with a torsion spring, the sum of thrust magnitudes produced at such equilibrium becomes precisely the vehicle’s weight whereas their directions are aligned vertically in the world frame, resulting in eliminating internal forces at any base pose in steady state. This is achieved with no overactuation, contrarily to state-of-the-art solutions. However, if no spring is added, the amount of internal forces can still be made arbitrarily small by carefully choosing the characteristics of the attached links.
Figure 7: Possible implementation of the proposed concept in Fig. 1. The vehicle can perform a large range attitude maneuver around any axis while keeping the propellers vertical thanks to the 2-DoF joints between the base and the propellers.
By proving that the zero dynamics, represented by the passive joints dynamics, is asymptotically stable, any base pose can be rendered asymptotically stable for the proposed MAV under a dynamic state feedback controller, as long as the extended system state trajectories remain in a certain neighborhood, which is constructed based on applying 1-step of the dynamic extension algorithm [23] followed by feedback linearization on the I/O map in a neighborhood of the desired equilibrium. The remaining virtual input is then obtained by implementing a backstepping strategy, taking the base configuration error as the left-trivialized error on $SE(3)$. Furthermore, to implement this dynamic controller efficiently, we introduced a method to evaluate the acceleration recursively without resorting to the inversion of the generalized mass matrix or numerical differentiation schemes.
While the simulation shows that the proposed MAV can perform decoupled motions even in presence of non-idealities, this might require the system trajectories to stay close to the neighborhood boundary, especially when the desired maneuver makes the actively-tilting thrust approach zero, leading to surges in the control variables. Future work will focus on developing a control scheme that overcomes this limitation in parallel with investigating the case with more generic (non-coplanar) propellers arrangements, relaxing the assumptions on the actively-tilting propeller link, and extending the control scheme to handle trajectory tracking problems. Providing experimental validation of the proposed concept will also receive attention. A possible mechanical design is displayed in Fig. 7, which supports the potential technological value of the proposed work; however, the realization of the mechatronics prototype requires considerable additional effort and represents non-negligible future work which is beyond the scope of the research presented in this paper.
Acknowledgments
The authors gratefully acknowledge Quentin Sablé (University of Twente; q.l.g.sable@utwente.nl) for producing some of the figures used in this work.
VII Proofs
VII-A Proof of Lemma 1
VII-A1 Necessity Proof
Assume that at least one propeller, labeled $l$, is not vertical in the inertial frame, hence $C^{0}_{l}(\bar{C}^{0}_{1},\alpha_{l})\,e_{3}\neq e_{3}$. The thrust force vector of propeller $i$, expressed in the inertial frame, is denoted by $T^{0}_{i}=f_{i}\,\alpha_{i}$. When the MAV is at equilibrium, the resultant thrust force in the inertial frame must satisfy $T^{0}_{tot}=m_{\rm tot}g\,(\,0\,0\,1\,)^{T}$ to compensate for the gravitational force. Thus, this constraint must hold at equilibrium
$\sum_{i=1}^{l-1}f_{i}\,(0\;0\;1)^{T}+f_{l}\,(\alpha_{lx}\;\alpha_{ly}\;\alpha_{lz})^{T}=m_{\rm tot}g\,(0\;0\;1)^{T}$
$\sqrt{\alpha_{lx}^{2}+\alpha_{ly}^{2}+\alpha_{lz}^{2}}=1$ (29)
Expanding this yields
${f}_{l}\,\alpha_{lx}=0,\,\,{f}_{l}\,\alpha_{ly}=0,\,\,{f}_{l}\,\alpha_{lz}=m_{\rm tot}g-\sum^{l-1}_{i=1}f_{i}$ (30)
If ${f}_{l}>0$ then (30) gives $(\,\alpha_{lx},\,\alpha_{ly}\,)=(0,0)$. This makes (29) have only two possible solution $\alpha_{lz}=\pm 1$. Setting $\alpha_{lz}=1$ in (30) yields $\sum^{l-1}_{i=1}f_{i}<m_{\rm tot}g$. This solution for $\alpha_{l}$ is in a contradiction with the assumption as it implies that $C^{0}_{l}(\bar{C}^{0}_{1},\alpha_{l})\,e_{3}=e_{3}$ is true, i.e. propeller $l$ is vertical in the inertial frame, and (30) admits a solution of $f_{i}=\frac{1}{l}m_{\rm tot}g$. As a result, $f\in\mathcal{F}_{\rm min}^{\rm eq}$ and no internal force is generated at equilibrium. If instead ${f}_{l}>0$ and $(\,\alpha_{lx},\,\alpha_{ly},\,\alpha_{lz}\,)=(0,0,-1)$, then from (30) we have $\sum^{l-1}_{i=1}f_{i}>m_{\rm tot}g$. While this is not a contradiction, it leads to $f\in\mathcal{F}_{\rm int.for.}^{\rm eq}$. Hence, the assumption $C^{0}_{l}(\bar{C}^{0}_{1},\alpha_{l})\,e_{3}\neq e_{3}$ does not hold when $f\in\mathcal{F}_{\rm min}^{\rm eq}$, proving the necessary condition by contradiction. By induction, the same steps can be easily repeated to cases where two or more propellers assume $C^{0}_{l}(\bar{C}^{0}_{1},\alpha_{l})\,e_{3}\neq e_{3}$.
VII-A2 Sufficiency Proof
If all propellers point up at equilibrium, i.e. $C^{0}_{i}(\bar{C}^{0}_{1},\alpha_{i})\,e_{3}=e_{3}$ $\forall i\in\{1,\dots,l\}$, then this renders all direction vectors $\alpha_{i}$’s equal to $(\,0\,0\,1\,)^{T}$, which automatically satisfies (28) with $f_{i}=\frac{1}{l}m_{\rm tot}g$. Hence, $f\in\mathcal{F}_{\rm min}^{\rm eq}$.
VII-B Proof of Proposition 1
The proof is constructive and divided into two parts. In the first part, Statement 1 is proven by finding PJD2 joint equilibrium configurations $q_{te}^{[2]}$ from (8) for any given base pose $C^{0}_{1e}\in SE(3)$ and showing that the associated equilibrium thrust $f_{e}^{[2]}$ belongs to the minimal equilibrium thrust $\mathcal{F}_{\rm min}^{\rm eq}$ defined in (9) and remains fixed for any such pose $C^{0}_{1e}$, rendering Statement 2 an immediate corollary as the condition $f_{e}^{[2]}\in\mathcal{F}_{\rm min}^{\rm eq}$ implies that no excess thrust is present at any $C^{0}_{1e}$, i.e. $f_{e}^{[2]}\notin\mathcal{F}_{\rm int.for.}^{\rm eq}$ (Definition 9).
In the second part, we start by solving (7) for $q_{te}^{[1]}$. We then illustrate that, when $\tau_{dj}\neq 0$, there is no joint equilibrium satisfying $C^{0}_{j}(C^{0}_{1e},q^{[1]}_{je})e_{3}=e_{3}$, hence by Lemma 1, no minimal thrust is produced at these euilibria $q^{[1]}_{je}$, which is the subject of statement 3. Once again, statement 4 follows as a result since $C^{0}_{j}(C^{0}_{1e},q^{[1]}_{je})e_{3}\neq e_{3}$ necessarily implies $f_{e}^{[1]}\in\mathcal{F}_{\rm int.for.}^{\rm eq}$. However, we observe that $f_{e}^{[1]}$, when $\tau_{dj}\neq 0$, $f_{e}^{[1]}$ can be made arbitrarily close to the set $\mathcal{F}_{\rm min}^{\rm eq}$ (9), i.e. $q_{te}^{[1]}$ is arbitrarily close to meet $C^{0}_{j}(C^{0}_{1e},q^{[1]}_{je})e_{3}=e_{3}$, by changing the link parameters.
VII-B1 Statement 1 Proof
Expanding the LHS of (8) or (7) for any $B_{j}$ $\forall j=2,3,4$ yields the pair of equations for $B_{j}$
$g_{t_{j}}=-S^{T}_{j}W_{grav,B_{j}}^{0}=(\,\,-S^{T}_{j1}\,\,-S^{T}_{j2}\,\,)^{T}\,M_{B_{j}}^{0}G^{0}$ (31)
where the instantaneous spatial joint screws are given by $S_{j1}(C^{0}_{1})=\operatorname{Ad}_{C^{0}_{1}}Y_{j1}$ and $S_{j2}(C^{0}_{1},q_{j1})=\operatorname{Ad}_{C^{0}_{j1}}\operatorname{Ad}^{-1}_{A^{0}_{j1}}Y_{j2}$, with $A^{0}_{j1}$ being the configuration of the intermediate link $\bar{B}_{j}$ frame, placed at the center of rotation of $q_{j1}$, connecting $B_{1}$ and $B_{j}$ when the vehicle is in the home configuration shown in Fig. 1, and $C^{0}_{j1}$ being its configuration in a generic vehicle pose. $Y_{ji}\in se(3)$ is the constant spatial screw of $q_{ji}$ in home configuration. For details on screw theory, consult [25].
Now, consider the LHS of first equation in (31), which is written as
$\begin{split}g_{t_{j1}}&=-S^{T}_{j1}M_{B_{j}}^{0}(\,\,0_{1\times 5}\,\,-{g}\,\,)^{T}\\ &=Y^{T}_{j1}\operatorname{Ad}^{T}_{C^{0}_{1}}\operatorname{Ad}^{-T}_{C^{0}_{j}}I^{j}_{B_{j}}\operatorname{Ad}^{-1}_{C^{0}_{j}}\bar{e}^{6}_{6}g\end{split}$ (32)
where $I^{j}_{B_{j}}$ is the diagonal $B_{j}$’s CoM-referred inertia tensor. $\bar{e}^{k}_{i}\in\mathbb{R}^{k}$ is a zero vector with $1$ in the $i$-th component. Since the adjoint matrix $\operatorname{Ad}^{-T}_{C^{0}_{j}}$ has the properties
$\begin{split}\operatorname{Ad}^{-T}_{C^{0}_{j}}&=\operatorname{Ad}^{T}_{C^{j}_{0}}=(\operatorname{Ad}_{C^{j}_{1}C^{1}_{0}})^{T}=(\operatorname{Ad}_{C^{j}_{1}}\operatorname{Ad}_{C^{1}_{0}})^{T}\\ &=\operatorname{Ad}^{T}_{C^{1}_{0}}\operatorname{Ad}^{T}_{C^{j}_{1}}=\operatorname{Ad}^{-T}_{C^{0}_{1}}\operatorname{Ad}^{-T}_{C^{1}_{j}}\end{split}$ (33)
substituting $\operatorname{Ad}^{-T}_{C^{0}_{j}}=\operatorname{Ad}^{-T}_{C^{0}_{1}}\operatorname{Ad}^{-T}_{C^{1}_{j}}$ in (32) gives
$g_{t_{j1}}=Y^{T}_{j1}\operatorname{Ad}^{-T}_{C^{1}_{j}}I^{j}_{B_{j}}\operatorname{Ad}^{-1}_{C^{0}_{j}}\bar{e}^{6}_{6}g=Y^{T}_{j1}\operatorname{Ad}^{-T}_{C^{1}_{j}}I^{j}_{B_{j}}\bar{r}_{j}g$ (34)
where $\bar{r}_{j}$ denotes
$\bar{r}_{j}=\operatorname{Ad}^{-1}_{C^{0}_{j}}\bar{e}^{6}_{6}$ (35)
Observe that substituting $\bar{r}_{j}=\bar{e}^{6}_{6}$ in (34) reduces it to
$\begin{split}g_{t_{j1}}&=Y^{T}_{j1}\operatorname{Ad}^{-T}_{C^{1}_{j}}I^{j}_{B_{j}}\bar{e}^{6}_{6}g\\ &=Y^{T}_{j1}\operatorname{Ad}^{-T}_{C^{1}_{j}}\bar{e}^{6}_{6}m_{p}g\end{split}$ (36)
where $g_{t_{j1}}$ now depends only on $q_{j}$. Since the constant screw is $Y^{T}_{j1}=(\,\,(\bar{e}^{3}_{1})^{T}\,\,\,\sigma_{j}\frac{a}{2}(\bar{e}^{3}_{3})^{T}\,\,)$, where $\sigma_{j}=1$ for $j=2,5$ and $\sigma_{j}=-1$ otherwise, (36) can be manipulated further as
[…]
where $C^{1}_{j}=\left(\begin{smallmatrix}R^{1}_{j}&P^{1}_{j}\\ 0_{1\times 3}&1\end{smallmatrix}\right)$. Hence setting $\bar{r}_{j}=\bar{e}^{6}_{6}$ satisfies the first equation in (8).
Next, let us consider the second equation in (8). Simplifying using the properties in (33), and substituting the expression of the constant screw $Y_{j2}$, configuration $A^{0}_{j1}$ and $\bar{r}_{j}$ (35) yield
$\begin{split}g_{t_{j2}}&=Y_{j2}^{T}\operatorname{Ad}^{-T}_{A^{0}_{j1}}\operatorname{Ad}^{T}_{C^{0}_{j1}}\operatorname{Ad}^{-T}_{C^{0}_{j}}I^{j}_{B_{j}}\operatorname{Ad}^{-1}_{C^{0}_{j}}\bar{e}^{6}_{6}g\\ &=Y_{j2}^{T}\operatorname{Ad}^{-T}_{A^{0}_{j1}}\operatorname{Ad}^{-T}_{C^{j1}_{j}}I^{j}_{B_{j}}\bar{r}_{j}g\\ &=(\,\,(\bar{e}^{3}_{2})^{T}\,\,0_{1\times 3}\,\,)\operatorname{Ad}^{-T}_{C^{j1}_{j}}I^{j}_{B_{j}}\bar{r}_{j}g\end{split}$ (38)
Setting $\bar{r}_{j}=\bar{e}^{6}_{6}$ gives
$g_{t_{j2}}$ $=(\,\,(\bar{e}^{3}_{2})^{T}\,\,0_{1\times 3}\,\,)\operatorname{Ad}^{-T}_{C^{j1}_{j}}\bar{e}^{6}_{6}m_{p}g$ (39)
$=(\,\,(\bar{e}^{3}_{2})^{T}\,\,0_{1\times 3}\,\,)\left(\begin{smallmatrix}-[P^{j1}_{j}]^{T}R^{j1}_{j}\bar{e}^{3}_{3}\\ R^{j1}_{j}\bar{e}^{3}_{3}\end{smallmatrix}\right)m_{p}g$
$=-(\bar{e}^{3}_{2})^{T}[P^{j1}_{j}]^{T}R^{j1}_{j}\bar{e}^{3}_{3}\,m_{p}g$
$=\,(\,\,\,-\cos(q_{j2e})\,\,\,0\,\,\,\sin(q_{j2e})\,\,\,)\left(\begin{smallmatrix}\sin(q_{j2e})\\ 0\\ \cos(q_{j2e})\end{smallmatrix}\right)c\,m_{p}g$
[…]
We conclude that $\bar{r}_{j}=\bar{e}^{6}_{6}$ satisfies (8). Observe that this condition implies that all propellers are aligned vertically in the inertial frame at equilibria for any $C^{0}_{1e}\in SE(3)$. To see this, expand (35) as
$\bar{r}_{j}$ $=\bar{e}^{6}_{6}=\operatorname{Ad}^{-1}_{C^{0}_{j}}\bar{e}^{6}_{6}$ (40)
$\left(\begin{smallmatrix}0_{3\times 1}\\ \bar{e}^{3}_{3}\end{smallmatrix}\right)$ $=\left(\begin{smallmatrix}R^{j}_{0}&0_{3\times 3}\\ -R^{j}_{0}[P^{0}_{j}]&R^{j}_{0}\end{smallmatrix}\right)\left(\begin{smallmatrix}0_{3\times 1}\\ \bar{e}^{3}_{3}\end{smallmatrix}\right)=\left(\begin{smallmatrix}0_{3\times 1}\\ R^{j}_{0}\bar{e}^{3}_{3}\end{smallmatrix}\right)$
Hence $\bar{e}^{3}_{3}=R^{j}_{0}(R^{0}_{1e},q^{[2]}_{je})\bar{e}^{3}_{3}\,\,\forall j=2,3,4,5$, or alternatively $C^{0}_{j}e_{3}=e_{3}$ using Lemma 1 notation. This also gives the expression of $q^{[2]}_{je}$ by solving $\bar{r}_{j}=\bar{e}^{6}_{6}$ for $q^{[2]}_{je}$ as
$\begin{split}R^{1e}_{0}\bar{e}^{3}_{3}=R^{1}_{j}(q^{[2]}_{je})\bar{e}^{3}_{3}=R_{x}(q^{[2]}_{j1e})R_{y}(q^{[2]}_{j2e})\bar{e}^{3}_{3}\end{split}$ (41)
Thus we have the solution $q^{[2]}_{je}$ as
$\begin{split}l_{s1}&=\frac{-r_{32}}{\sqrt{1-r_{31}^{2}}},\,l_{c1}=\sqrt{1-l_{s1}^{2}}\\ l_{s2}&=r_{31},\,l_{c2}=\sqrt{1-l_{s2}^{2}}\\ q^{[2]}_{jie}&=\operatorname{atan}(l_{si},l_{ci})\,\,\forall i=1,2,\,j=2,3,4,5\end{split}$ (42)
where $r_{mn}$ is the element $(m,n)$ in any desired base orientation at PJD2 equilibrium $R^{0}_{1e}$. Notice that $q^{[2]}_{je}$ is equal for any attached link $B_{j}$. Moreover, it is always possible to obtain this solution $q^{[2]}_{je}$ (42) for every $R^{0}_{1}\in SO(3)$ since aligning the unit vector $R^{1}_{j}(q^{[2]}_{je})\bar{e}^{3}_{3}=:z^{1}_{j}$, where $z^{1}_{j}\in\mathbb{S}^{2}$ is the z-axis of $\mathcal{F}_{p_{j}}$ expressed in $\mathcal{F}_{b}$, with any desired unit vector $z^{1}_{0}\in\mathbb{S}^{2}$ in $\mathcal{F}_{b}$, where $R^{1}_{0}\bar{e}^{3}_{3}=:z^{1}_{0}$, requires only 2 variables, i.e. $q^{[2]}_{j1e},q^{[2]}_{j2e}$, as $\operatorname{dim}(\mathbb{S}^{2})=2$ when the manifold $\mathbb{S}^{2}$ is embedded in the 3-dimensional Euclidean vector space.
Now, we show that the associated thrust $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ and fixed for any $C^{0}_{1e}$. The dynamics of base at any static equilibrium reduces to
$\begin{split}g_{b}(C^{0}_{1e},q_{te})=&-W^{0}_{app,b}(f_{e},q_{te},q_{5e})\end{split}$ (43)
which reads explicitly as
$\sum_{j=2}^{5}\operatorname{Ad}^{T}_{(C^{0}_{f_{j}})^{-1}}\bar{f}_{j}f_{je}=\sum_{j=1}^{5}\operatorname{Ad}^{-T}_{(C^{0}_{j})}I^{j}_{B_{j}}\operatorname{Ad}^{-1}_{(C^{0}_{j})}\bar{e}^{6}_{6}g$ (44)
Substituting $\bar{r}_{j}=(\,\,0_{3\times 1}\,\,\hat{r}_{j}\,\,)$ from (35) yields
$\begin{split}\sum_{j=2}^{5}\begin{pmatrix}(Ik_{j}-[P^{1}_{j}]^{T})R^{1}_{j}\bar{e}^{3}_{3}\\ R^{1}_{j}\bar{e}^{3}_{3}\end{pmatrix}f_{je}&=g\,m_{b}\begin{pmatrix}0_{3\times 1}\\ R^{1}_{0}\bar{e}^{3}_{3}\end{pmatrix}\\ &\quad+g\sum_{j=2}^{5}m_{j}\begin{pmatrix}-[P^{1}_{j}]^{T}R^{1}_{j}\hat{r}_{j}\\ R^{1}_{j}\hat{r}_{j}\end{pmatrix}\end{split}$
where $I$ is the identity matrix. Using the property $R^{1}_{j}[R^{j}_{1}P^{1}_{j}]=-[P^{1}_{j}]^{T}R^{1}_{j}$ and $R^{1}_{j}\hat{r}_{j}=R^{1}_{0}\bar{e}^{3}_{3}$, we get
$\begin{split}&\sum_{j=2}^{5}R^{1}_{j}(Ik_{j}+[R^{j}_{1}P^{1}_{j}])\bar{e}^{3}_{3}f_{je}=g\Big(\sum_{j=2}^{5}R^{1}_{j}[R^{j}_{1}P^{1}_{j}]\hat{r}_{j}m_{j}\Big)\\ &\sum_{j=2}^{5}R^{1}_{j}\bar{e}^{3}_{3}f_{je}=g\sum_{j=2}^{5}R^{1}_{j}\hat{r}_{j}\Big(\frac{1}{4}m_{b}+m_{j}\Big)\end{split}$
Since in PJD2, $q^{[2]}_{je}\,\forall j=2,3,4,5$ are equal, $R^{1}_{j}$ can be taken as a common factor, hence pre-multiplying by $R^{j}_{1}$ gives
$\begin{split}&\sum_{j=2}^{5}k_{j}f^{[2]}_{je}\bar{e}^{3}_{3}+\sum_{j=2}^{5}[R^{j}_{1}P^{1}_{j}]f^{[2]}_{je}\bar{e}^{3}_{3}=g\Big(\sum_{j=2}^{5}[R^{j}_{1}P^{1}_{j}]m_{j}\Big)\bar{e}^{3}_{3}\\ &\bar{e}^{3}_{3}\sum_{j=2}^{5}f^{[2]}_{je}=\bar{e}^{3}_{3}\sum_{j=2}^{5}g\Big(\frac{1}{4}m_{b}+m_{j}\Big)\end{split}$
After $\hat{r}_{j}=\bar{e}^{3}_{3}$ is substituted. The second equation is satisfied with $f^{[2]}_{je}={g(\frac{1}{4}m_{b}+m_{j})}$, hence $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$. Substituting this $f^{[2]}_{je}$ make the first equation reduce to
$\sum_{j=2}^{5}k_{j}{g(\frac{1}{4}m_{b}+m_{j})}\bar{e}^{3}_{3}+\sum_{j=2}^{5}[R^{j}_{1}P^{1}_{j}]\bar{e}^{3}_{3}=0$ (48)
thus $\sum_{j=2}^{5}k_{j}{g(\frac{1}{4}m_{b}+m_{j})}=0$ since $\sum_{j=2}^{5}[R^{j}_{1}P^{1}_{j}]\bar{e}^{3}_{3}=0$ is always true at any $q^{[2]}_{je}$ (42). We conclude that the associated thrust at $q^{[2]}_{je}$ is, when $m_{j}=m_{p}\,\forall j=2,3,4,5$
$f^{[2]}_{je}={g(\frac{1}{4}m_{b}+m_{p})},\text{ with $k_{i}$ chosen by}\sum_{j=2}^{5}k_{i}=0$ (49)
VII-B2 Statement 2 Proof
Since $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ is proven in Statement 1, it follows from Definition 1 that no internal forces are generated $f^{[2]}_{e}\notin\mathcal{F}_{\rm int.for.}^{\rm eq}$ at the corresponding equilibrium configurations associated with this $f^{[2]}_{e}$, i.e. $q^{[2]}_{te}$ in (42) and any desired $C^{0}_{1e}\in SE(3)$.
VII-B3 Statement 3 Proof
In Sec. VII-B1, it is found that when all propellers point up in the inertial frame, i.e. $\bar{r}_{j}=\bar{e}^{6}_{6}\text{ or }C^{0}_{j}e_{3}=e_{3}\,\,\forall j=2,3,4,5$, the LHS of (7) becomes $g_{t_{j}}(C^{0}_{1e},q_{te})=0$. Since the RHS of of it for any $j$ is equal to $\left(\begin{smallmatrix}\tau_{dj}\\ 0\end{smallmatrix}\right)$, which is not zero when $\tau_{dj}\neq 0$, $C^{0}_{j}(C^{0}_{1e},q^{[1]}_{je})e_{3}\neq e_{3}$ must hold true. Thus, from Lemma 1, no minimal thrust is produced at any PJD1 equilibrium $f^{[1]}_{e}\notin\mathcal{F}_{\rm min}^{\rm eq}\,\forall\tau_{dj}\neq 0$.
For completeness, we proceed to find conditions on $q^{[1]}_{je}$, similar to $\bar{r}_{j}=\bar{e}^{6}_{6}$ obtained for PJD2, that must be met at any base equilibrium $C^{0}_{1e}$. We start from (38) and substitute the values of the constant screws, constant transformations and $\bar{r}_{j}=(\,\,0_{3\times 1}\,\,\hat{r}_{j}\,\,)$, which yields
$(\bar{e}^{3}_{2})^{T}[P^{j1}_{j}]^{T}R^{j1}_{j}\hat{r}_{j}m_{j}g=0$ (50)
which is satisfied when $\hat{r}_{j}=(\,\,0\,\,\hat{r}_{j,1}\,\,\hat{r}_{j,2}\,\,),\,\hat{r}_{j,2}=\sqrt{1-\hat{r}_{j,1}^{2}}$. in PJD1, (34) has the RHS equal to $S^{T}_{j1}W_{app,B_{j}}^{0}$. This can be expressed as
$\begin{split}(\sigma_{j}\frac{a}{2}(\bar{e}^{3}_{3})^{T}-(\bar{e}^{3}_{1})^{T}[P^{1}_{j}]^{T})&R^{1}_{j}\big(\hat{r}_{j}m_{j}g-\bar{e}^{3}_{3}f^{[1]}_{je}\big)\\ &+(\bar{e}^{3}_{1})^{T}R^{1}_{j}\bar{e}^{3}_{3}k_{j}f^{[1]}_{je}=0\end{split}$ (51)
since
$Y^{T}_{j1}\operatorname{Ad}^{-T}_{C^{1}_{j}}=\left(\begin{smallmatrix}\cos(q_{j2e})&0&\sin(q{j2e})&0&c\cos(q{j2e})&0\end{smallmatrix}\right)$ (52)
hence $q^{[1]}_{j2e}$ is constraint by
$\cos(q^{[1]}_{j2e})=\pm\sqrt{\frac{k_{j}^{2}(f^{[1]}_{je})^{2}}{c^{2}m_{j}^{2}g^{2}\hat{r}_{j,1}^{2}+k_{j}^{2}(f^{[1]}_{je})^{2}}}$ (53)
Lastly, from (46), we have these conditions on $R^{1}_{j}(q^{[1]}_{je}),\,f^{[1]}_{je}$
$\begin{split}&\sum_{j=2}^{5}R^{1}_{j}k_{j}\bar{e}^{3}_{3}f^{[1]}_{je}=0,\,\,R^{1}_{j}\hat{r}_{j}=R^{1}_{0}\bar{e}^{3}_{3}\\ &\sum_{j=2}^{5}[P^{1}_{j}](R^{1}_{j}\bar{e}^{3}_{3}f^{[1]}_{je}-R^{1}_{j}\hat{r}_{j}gm_{j})=0\\ &\sum_{j=2}^{5}R^{1}_{j}\bar{e}^{3}_{3}f^{[1]}_{je}-\sum_{j=2}^{5}R^{1}_{j}\hat{r}_{j}g\Big(\frac{1}{4}m_{b}+m_{j}\Big)=0\end{split}$ (54)
VII-B4 Statement 4 Proof
By definition 1, and Statement 3, $f^{[1]}_{e}\in\mathcal{F}_{\rm int.for.}^{\rm eq}$ immediately follows as a corollary. The amount of internal force can be made arbitrarily small if solutions of (8) and (7) are made arbitrarily close, which is realized by making the term $\frac{k_{j}}{cm_{j}}=0$, i.e. reducing the drag-to-thrust coefficient or increasing the pendulum length or mass.
VII-C Proof of Theorem 1
The proof relies on the relative degree notion for MIMO affine nonlinear systems [26] and the dynamic extension algorithm for invertible systems [23] to prove Statement 1. Afterwards, we show that Statement 2 is true by a standard Lyapunov argument applied to the zero dynamics in the normal form.
VII-C1 Statement 1 Proof
To compute the relative degree of $\Sigma_{2}$ w.r.t. $x_{21}$, we take time derivatives of the output $x_{21}$
$\begin{split}\dot{h}_{y}=\dot{V}^{0}_{1}\end{split}$ (55)
From (3) $\dot{V}^{0}_{1}$ is expressed as
$M_{bb}\dot{V}^{0}_{1}=-M_{bt}\ddot{q}_{t}-h_{b}-g_{b}+W^{0}_{app,b}$ (56)
Following the approach of [27] we can express $\ddot{q}_{t}$ in terms of $\dot{V}^{0}_{1}$, $W^{0}_{app,t}$ and ${\tau}_{f}$, hence
$\begin{split}\ddot{q}_{t}=&C_{t1}-M_{tt}^{-1}M_{bt}^{T}\dot{V}^{0}_{1}+M_{tt}^{-1}W_{app,t}^{0}\end{split}$ (57)
where $C_{t1}$ is defined as
$C_{t1}:=M_{tt}^{-1}(\tau_{f}-h_{t}-g_{t})$ (58)
Substituting in (56) gives
$\begin{split}(M_{bb}-M_{bt}M_{tt}^{-1}M_{bt}^{T})\dot{V}^{0}_{1}=&-M_{bt}C_{t1}-h_{b}-g_{b}\\ &+W^{0}_{app,b}-M_{bt}M_{tt}^{-1}W_{app,t}^{0}\end{split}$ (59)
which can be rewritten as
$\dot{h}_{y}=\dot{V}^{0}_{1}=F_{1}+M_{b}^{-1}D_{1}u$ (60)
where $M_{b}$, $D_{1}$ and $F_{1}$ are defined as, with $j=2,3,4$
$\begin{split}M_{b}=&(M_{bb}-M_{bt}M_{tt}^{-1}M_{bt}^{T})\\ B_{j}=&\big(I_{6}-M^{0}_{B_{j}}{S}_{j}(S_{j}^{T}M^{0}_{B_{j}}{S}_{j})^{-1}S_{j}^{T}\big)\\ \bar{D}_{1}=&\left(\begin{array}[]{ccc}B_{j}\operatorname{Ad}^{T}_{(C^{0}_{f_{j}})^{-1}}\bar{f}_{j}&\cdots&\operatorname{Ad}^{T}_{(C^{0}_{f_{5}})^{-1}}\bar{f}_{5}\end{array}\right)\\ D_{1}=&\left(\begin{array}[]{cccc}\bar{D}_{1}&0_{6\times 2}\end{array}\right)\\ F_{1}=&M_{b}^{-1}(-M_{bt}C_{t1}-h_{b}-g_{b})\end{split}$ (61)
Note that $(S_{j}^{T}M^{0}_{B_{j}}{S}_{j})\in R^{2\times 2}$ is a symmetric positive definite matrix.
Since the decoupling matrix $M_{b}^{-1}D_{1}$ is singular due to $D_{1}$ being singular of rank 4 $\forall{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}$, system (15) does not have a vector relative degree in any open neighbourhood of the equilibrium ${x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}$. Thus, the system is not I/O feedback linearizable by any static state feedback as the relative degree is invariant under this kind of feedback  [26].
However, extending the system dynamics $\Sigma_{2}$ with an input precompensator, by applying the version of the dynamic extension algorithm presented in  [23] for invertible nonlinear systems, may result in a system for which the vector relative degree is well-defined. Assume the precompensator takes this pure integrator form (16). Let $\bar{x}:=(x_{12},x_{2},q_{5},z)^{T}\in\bar{\mathbb{X}}$ be the extended state and $w:=(\bar{w},u_{v_{1}},u_{v_{2}})^{T}$ the new input. The extended system $\Sigma_{\bar{2}}$ can be written as $\Sigma_{\bar{2}}:\dot{\bar{x}}=\bar{F}+\bar{G}w$ with $\bar{F}:=F+G\bar{u}$, where $\bar{u}=(z,0,0)^{T}$ and $\bar{G}$ is given by
$\bar{G}=\left(\begin{array}[]{c}0_{18\times 6}\\ 0_{2\times 4}\quad I_{2\times 2}\\ I_{4\times 4}\quad 0_{2\times 2}\end{array}\right)$ (62)
Now, we can proceed with computing the vector relative degree of $\Sigma_{\bar{2}}$ w.r.t $x_{21}$ by differentiating $\dot{h}_{y}$ w.r.t time
$M_{bb}\ddot{V}^{0}_{1}=-\dot{M}_{bt}\ddot{q}_{t}-{M}_{bt}\dddot{q}_{t}+\dot{W}^{0}_{app,b}-\dot{M}_{bb}\dot{V}^{0}_{1}-\dot{h}_{b}-\dot{g}_{b}$
by substituting this expression for $\dddot{q}_{t}$
$\dddot{q}_{t}=C_{t2}-M_{tt}^{-1}M_{bt}^{T}\ddot{V}^{0}_{1}+M_{tt}^{-1}\dot{W}_{app,t}^{0}$ (63)
where $C_{t2}$ is defined as
$C_{t2}:=M_{tt}^{-1}\big((\dot{\tau}_{f}-\dot{h}_{t}-\dot{g}_{t})-(\dot{M}_{bt}^{T}\dot{V}^{0}_{1}+\dot{M}_{tt}\ddot{q}_{t})\big)$ (64)
and after some algebraic manipulation it yields
$\begin{split}M_{b}\ddot{V}^{0}_{1}=&-\dot{M}_{bt}(C_{t1}+M_{tt}^{-1}W_{app,t}^{0})-{M}_{bt}C_{t2}\\ &+(\dot{M}_{bt}M_{tt}^{-1}M_{bt}^{T}-\dot{M}_{bb})\dot{V}^{0}_{1}-\dot{h}_{b}-\dot{g}_{b}\\ &-{M}_{bt}M_{tt}^{-1}\dot{W}_{app,t}^{0}+\dot{W}^{0}_{app,b}\end{split}$ (65)
where the $\dot{h}_{b},\dot{g}_{b},\dot{h}_{t},\dot{g}_{t},\ddot{M},\ddot{S}_{j},\dot{W}^{0}$ are given by, $j=2,3,4$
$\begin{split}\dot{h}_{b}=&\sum^{4}_{j=2}\big((M^{0}_{B_{j}}\ddot{S}_{j}+\ddot{M}^{0}_{B_{j}}{S}_{j}+2\dot{M}^{0}_{B_{j}}\dot{S}_{j})\dot{q}_{j}\\ &+(M^{0}_{B_{j}}\dot{S}_{j}+\dot{M}^{0}_{B_{j}}{S}_{j})\ddot{q}_{j}\big)+\ddot{M}_{bb}{V}^{0}_{1}+\dot{M}_{bb}\dot{V}^{0}_{1}\\ \dot{h}_{t}=&\operatorname{col}\big(\dot{S}_{j}^{T}((M^{0}_{B_{j}}\dot{S}_{j}+\dot{M}^{0}_{B_{j}}S_{j})\dot{q}_{j}+\dot{M}^{0}_{B_{j}}{V}^{0}_{1})\\ &\quad\quad+S_{j}^{T}((M^{0}_{B_{j}}\ddot{S}_{j}+\ddot{M}^{0}_{B_{j}}S_{j}+2\dot{M}^{0}_{B_{j}}\dot{S}_{j})\dot{q}_{j}\\ &\quad\quad+(M^{0}_{B_{j}}\dot{S}_{j}+\dot{M}^{0}_{B_{j}}S_{j})\ddot{q}_{j}+\ddot{M}^{0}_{B_{j}}{V}^{0}_{1}+\dot{M}^{0}_{B_{j}}\dot{V}^{0}_{1})\big)\\ \dot{g}_{b}&=-(\dot{M}_{bb}+\dot{M}^{0}_{B_{5}})G^{0}\\ \dot{g}_{t}&=-\dot{S}_{t}^{T}\operatorname{col}({M}^{0}_{B_{j}}G^{0})-S_{t}^{T}\operatorname{col}(\dot{M}^{0}_{B_{j}}G^{0})\\ \ddot{M}^{0}_{k}&=M^{0}_{k}\operatorname{ad}^{2}_{{V}^{0}_{k}}+(\operatorname{ad}^{2}_{{V}^{0}_{k}})^{T}M^{0}_{k}+2\operatorname{ad}^{T}_{{V}^{0}_{k}}M^{0}_{k}\operatorname{ad}_{V^{0}_{k}}\\ &-M^{0}_{k}\operatorname{ad}_{\dot{V}^{0}_{k}}-\operatorname{ad}^{T}_{\dot{V}^{0}_{k}}M^{0}_{k},\text{ if }k=5\text{ then }{V}^{0}_{k}={V}^{0}_{1},\,\dot{V}^{0}_{k}=\dot{V}^{0}_{1}\\ \ddot{S}_{k}&=(\operatorname{ad}_{\dot{V}^{0}_{k}}+\operatorname{ad}^{2}_{V^{0}_{k}}){S}_{k},\,k=2,3,4\\ \dot{W}_{k}^{0}&=-\operatorname{ad}^{T}_{V_{k}^{0}}{W}_{k}^{0}+\operatorname{Ad}^{T}_{(C^{0}_{f_{5}})^{-1}}\dot{W}_{k}^{b},\,k=2,3,4,5\end{split}$
Note that ${M}^{0}_{B_{5}}$ depends only on $C^{0}_{1}$, i.e. ${M}^{0}_{B_{5}}(C^{0}_{1})$, due to the assumptions on link $B_{5}$. Hence, the input $\dot{q}_{5}$ does not appear in $\dot{M}^{0}_{B_{5}}(C^{0}_{1})$. Observe that (65) can be put in the form,
$\ddot{h}_{y}=\ddot{V}^{0}_{1}=F_{2}+M_{b}^{-1}D_{2}w$ (67)
where $F_{2}$ and $D_{2}$ are calculated from, with $j=2,3,4$
$\begin{aligned} D_{2}&=\left(\begin{array}[]{cccc}B_{j}\operatorname{Ad}^{T}_{(C^{0}_{f_{j}})^{-1}}\bar{f}_{j}&\cdots&\operatorname{Ad}^{T}_{(C^{0}_{f_{5}})^{-1}}\bar{f}_{5}&\bar{D}_{2}\end{array}\right)\\ \bar{D}_{2}&=\left(\begin{array}[]{cc}-\operatorname{ad}^{T}_{S_{51}}\operatorname{Ad}^{T}_{(C^{0}_{f_{5}})^{-1}}\bar{f}_{5}z_{5}&-\operatorname{ad}^{T}_{S_{52}}\operatorname{Ad}^{T}_{(C^{0}_{f_{5}})^{-1}}\bar{f}_{5}z_{5}\end{array}\right)\\ F_{2}&=M_{b}^{-1}\big(-\dot{M}_{bt}(C_{t1}+M_{tt}^{-1}W_{app,t}^{0})-{M}_{bt}C_{t2}\\ &\quad\quad\quad\quad+(\dot{M}_{bt}M_{tt}^{-1}M_{bt}^{T}-\dot{M}_{bb})\dot{V}^{0}_{1}-\dot{h}_{b}-\dot{g}_{b}+\bar{W}\big)\\ \bar{W}&=-\sum^{4}_{j=2}B_{j}\operatorname{ad}^{T}_{V_{j2}^{0}}W_{app,B_{j}}^{0}-\operatorname{ad}^{T}_{V_{1}^{0}}W_{app,B_{5}}^{0}\end{aligned}$
where $C_{t1}$, $C_{t2}$, $M_{b}$ and $B_{j}$ are given in (58), (64) and (61).
Because of the decoupling matrix $M_{b}^{-1}D_{2}$ of the extended system $\Sigma_{\bar{2}}$ being full rank equal to 6 in the open neighbourhood $\mathbb{U}_{\bar{x}_{e}}\subseteq\bar{\mathbb{X}}$ of equilibrium $\bar{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}\cup\mathbb{D}^{[i]}_{f_{e}}$, system $\Sigma_{\bar{2}}$ has a vector relative degree $(2,2,2,2,2,2)$. Hence it can be transformed to a normal form $\Sigma^{H}_{\tilde{2}}$ where the input/output map is linear by the means of the regular static state feedback w{w} in (17) and the local diffomorphism $H(\bar{x}):\mathbb{U}_{\bar{x}_{e}}\rightarrow\mathbb{U}_{\tilde{x}_{e}}$
$\begin{split}\left(\begin{array}[]{c}\zeta_{1}\\ \zeta_{2}\\ \eta\end{array}\right)&=H(\bar{x}):=\left(\begin{array}[]{c}h_{y}(x)\\ \dot{h}_{y}(\bar{x})\\ \Phi(\bar{x})\end{array}\right)\end{split}$ (69)
where $\Phi(\bar{x}):\mathbb{U}_{\bar{x}_{e}}\rightarrow\mathbb{T}^{6}\times\mathbb{R}^{6}$ is a sufficiently smooth function chosen such that, in the open neighbourhood $\mathbb{U}_{\bar{x}_{e}}$, $H(\bar{x})$ is a diffomorphism and $\frac{\partial\Phi}{\partial\bar{x}}\bar{G}=0$ holds. A possible choice of $\Phi$ is the states of the passive joints, as suggested in [27], $\Phi:=(q_{t},\dot{q}_{t})^{T}$.
Hence the internal dynamics can be expressed as
$\begin{split}\dot{\eta}_{1}&={\eta}_{2}\\ \dot{\eta}_{2}&=f_{\eta}({\eta}_{1},{\eta}_{2},{\zeta}_{1},\zeta_{2},\bar{z},C^{0}_{1})\\ &:=C_{t1}-M_{tt}^{-1}M_{bt}^{T}\zeta_{2}+M_{tt}^{-1}D_{3}\bar{z}\end{split}$ (70)
where $\eta$-dynamics represent the internal dynamics and $D_{3}$, $\bar{z}$ are calculated from
$\begin{split}\bar{z}&=\left(\begin{array}[]{ccc}z_{2}&z_{3}&z_{4}\end{array}\right)^{T}\\ D_{3}&=\operatorname{blkdiag}\big((k_{2}s_{22},0)^{T},(k_{3}s_{32},0)^{T},(k_{4}s_{42},0)^{T}\big)\end{split}$ (71)
Furthermore, we can map the controller states to the new coordinates $(z,q_{5})^{T}=f_{c}({\zeta},\eta,C^{0}_{1})$ where $f_{c}$ is a solution of the algebraic equations
$E=\tilde{D}\bar{z}+Cz_{5}$ (72)
where $\tilde{D},E,C$ are given by
[…]
Let $\bar{z}=\bar{f}_{c}$ where $\bar{f}_{c}$ is the first 3 elements in the map ${f}_{c}$. Hence, substituting back into $\eta$-dynamics in (70) yields the normal form $\Sigma^{H}_{\tilde{2}}(\zeta,\eta;C^{0}_{1},v)$
$\Sigma^{H}_{\tilde{2}}:\left\{\begin{split}\dot{\zeta}_{1}&=\zeta_{2}\\ \dot{\zeta}_{2}&=v\\ \dot{\eta}_{1}&={\eta}_{2}\\ \dot{\eta}_{2}&=f_{\eta}({\eta}_{1},{\eta}_{2},{\zeta}_{1},\zeta_{2},C^{0}_{1},\bar{f}_{c}({\zeta},\eta,C^{0}_{1}))\\ &:=C_{t1}-M_{tt}^{-1}M_{bt}^{T}\zeta_{2}+M_{tt}^{-1}D_{3}\bar{f}_{c}({\zeta},\eta,C^{0}_{1})\\ h_{y}&={\zeta}_{1}\end{split}\right$. (74)
This proves Statement 1.
VII-C2 Statement 2 Proof
$\Sigma_{3}$ in the new coordinates becomes $\Sigma^{H}_{3}(C^{0}_{1};\zeta_{1})$
$\dot{C}^{0}_{1}=[\zeta_{1}]{C}^{0}_{1}$ (75)
Thus, $\Sigma^{H}_{\tilde{2}}$ (74) interconnected with $\Sigma_{3}$ (75) is equivalent to the system (4) with a linear relationship between virtual input $v$ and output from the base spatial twist $V^{0}_{1}$.
The zero dynamics of the network ($\Sigma^{H}_{\tilde{2}}$, $\Sigma^{H}_{3}$) are the internal dynamics when $h_{y}(t)\equiv 0\,\forall t\geq 0$. This constraint is imposed for all times by setting the virtual input $v(t)=0\,\forall t\geq 0$ and the initial conditions of $\zeta$-dynamics to be $(\zeta_{1}(0),\zeta_{2}(0))=(0,0)$. As a result, the linear differential equations of $\zeta$ admit the solution $(\zeta_{1}(t),\zeta_{2}(t))=(0,0)\,\forall t\geq 0$. This also leads to $\Sigma^{H}_{3}$ (75) having a solution ${C}^{0}_{1}(t)={C}^{0}_{1}(0)\,\forall t\geq 0$. Thus, the zero dynamics manifold $\mathbb{Z}$ is then obtained as
$\begin{split}\mathbb{Z}&=\Bigg\{{\eta}(t)\in\mathbb{T}^{6}\times\mathbb{R}^{6},\,t\in\mathbb{R}_{\geq 0}\,\Bigr\rvert\dot{\eta}(t)=\left(\begin{array}[]{c}{\eta_{2}}(t)\\ f_{\eta}\big({\eta}_{1}(t),{\eta}_{2}(t),0,0,{C}^{0}_{1}(0),\bar{f}_{c}(0,\eta(t),{C}^{0}_{1}(0))\big)\end{array}\right)\Bigg\}\end{split}$
Remark 2.
In the original coordinates of system $(\Sigma_{\bar{2}},\Sigma_{3})$, one can obtain the zero dynamics manifold by transforming back the the initial conditions and the input constraint $v=0$ from the normal into the $\bar{x}$ coordinates. Hence, $\zeta_{1}(0)=x_{21}(0)=0$ while the constraint $\zeta_{2}(0)=\dot{x}_{21}(0)=0$ restricts the initial conditions of the dynamic extension states $(z(0),q_{5}(0))^{T}$ to be a solution of (60) with $\dot{V}^{0}_{1}(0)=0$. Let us call this solution $f_{c_{x}}$, then the zero dynamics manifold in the original coordinates $\mathbb{Z}_{x}$ is defined by
$\begin{split}\mathbb{Z}_{x}&=\Bigg\{\bar{x}(t)\in\bar{\mathbb{X}},\,t\in\mathbb{R}_{\geq 0}\,\Bigr\rvert\Sigma_{\bar{2}}(\bar{x};\tilde{w},x_{11}(0)),\,\tilde{w}=-D_{2}^{-1}M_{b}F_{2},\\ &\quad x_{11}(0)\in SE(3),x_{21}(0)=0,(z(0),q_{5}(0))^{T}=f_{c_{x}}\Bigg\}\end{split}$
What is remaining to complete the proof is to study the stability of the zero dynamics. When $\eta\in\mathbb{Z}$ we get,
[…]
The superscript $\tilde{(.)}$ donates the value of the corresponding variable evaluated at $({C}^{0}_{1}(t),\zeta_{1}(t),\zeta_{2}(t))=({C}^{0}_{1}(0),0,0)$. $\tilde{C}$ is the Coriolis matrix which is linear in passive joint velocities ${\eta}_{2}$ and given by
$\begin{split}\tilde{C}=&\operatorname{blockdiag}\big(\tilde{C}_{2}(\eta_{1_{2}},\eta_{2_{2}}),\tilde{C}_{3}(\eta_{1_{3}},\eta_{2_{3}}),\tilde{C}_{4}(\eta_{1_{4}},\eta_{2_{4}})\big)\\ \tilde{C}_{j}=&\left(\begin{array}[]{cc}S_{j1}^{T}M^{0}_{B_{j}}\dot{S}_{j1}-S_{j1}^{T}\operatorname{ad}_{S_{j2}}^{T}M^{0}_{B_{j}}S_{j1}\eta_{2_{j2}}&S_{j1}^{T}M^{0}_{B_{j}}\dot{S}_{j2}-S_{j1}^{T}\operatorname{ad}_{S_{j2}}^{T}M^{0}_{B_{j}}S_{j2}\eta_{2_{j2}}\\ S_{j2}^{T}M^{0}_{B_{j}}\dot{S}_{j1}-S_{j2}^{T}\operatorname{ad}_{S_{j1}}^{T}M^{0}_{B_{j}}S_{j1}\eta_{2_{j1}}&S_{j2}^{T}M^{0}_{B_{j}}\dot{S}_{j2}-S_{j2}^{T}\operatorname{ad}_{S_{j1}}^{T}M^{0}_{B_{j}}S_{j2}\eta_{2_{j1}}\end{array}\right)\end{split}$
where $j=2,3,4$.
A suitable choice of positive definite smooth Lyapunov candidate is the total energy of the rigid bodies connected by the passive joints comprising the internal dynamics $V(\eta):\mathbb{Z}\rightarrow\mathbb{R}_{>0}\,\forall\eta\neq\eta_{e},\,V(\eta_{e})=0$ $\forall\eta_{e}\in\mathbb{D}^{[i]}_{x_{e}}$
$V=\frac{1}{2}\eta_{2}^{T}\tilde{M}_{tt}(\eta_{1}-\eta_{1e})\eta_{2}+\tilde{U}(\eta_{1}-\eta_{1e})$ (80)
where $\tilde{U}(\eta_{1}-\eta_{1e})$ is the potential energy of the passive dynamics in the manifold $\eta\in\mathbb{Z}$ whose time derivative is
$\dot{\tilde{U}}$ $=\eta_{2}^{T}\frac{\partial\tilde{U}}{\partial(\eta_{1}-\eta_{1e})}$ (81)
$=-\eta_{2}^{T}{\tilde{S}_{t}^{T}(\eta_{1}-\eta_{1e})}\operatorname{col}({\tilde{M}_{B_{j}}(\eta_{1}-\eta_{1e})}G^{0})=\eta_{2}^{T}\tilde{g}_{t}$
where $G^{0}=\operatorname{col}(0_{5},-9.81\,m/s^{2})$. Taking the time derivatives of $V$ yields
$\begin{split}\dot{V}&=\eta_{2}^{T}\tilde{M}_{tt}\dot{\eta}_{2}+\frac{1}{2}\eta_{2}^{T}\dot{\tilde{M}}_{tt}{\eta}_{2}+\dot{\tilde{U}}\\ &=\eta_{2}^{T}(\frac{1}{2}\dot{\tilde{M}}_{tt}-\tilde{C}){\eta}_{2}\\ &-\eta_{2}^{T}D_{f}{\eta}_{2}-\eta_{2}^{T}\tilde{g}_{t}+\eta_{2}^{T}D_{3}\tilde{z}+\eta_{2}^{T}\tilde{g}_{t}\end{split}$ (82)
Since $\frac{1}{2}\dot{\tilde{M}}_{tt}-\tilde{C}$, with $\dot{\tilde{M}}_{tt}$ and $\tilde{C}$ respectively given in Table I and (79), is skew-symmetric, $\dot{V}$ reduces to
$\begin{split}\dot{V}&=-\eta_{2}^{T}D_{f}{\eta}_{2}+\eta_{2}^{T}D_{3}\tilde{z}\end{split}$ (83)
Thus for asymptotic stability of $\eta_{e}$, $\tilde{z}$ must satisfy
$\begin{split}\eta_{2}^{T}D_{3}\tilde{z}<\eta_{2}^{T}D_{f}{\eta}_{2}\end{split}$ (84)
Since the quadratic form $\eta_{2}^{T}D_{f}{\eta}_{2}$ can be bounded by
$\lambda_{\text{min}}(D_{f})\|\eta_{2}\|^{2}\leq\eta_{2}^{T}D_{f}\eta_{2}\leq\lambda_{\text{max}}(D_{f})\|\eta_{2}\|^{2}$
Applying Cauchy–Schwarz inequality to bound $|\eta_{2}^{T}D_{3}\tilde{z}|\leq\|\eta_{2}\|\,\|D_{3}\tilde{z}\|$ yields the bounds on $\tilde{z}$ $\forall{\eta}\in\mathbb{Z}-\{\eta_{e}\}$ to be (19) in original coordinates, where also this norm property $\|D_{3}\tilde{z}\|\leq\|D_{3}\|\|\tilde{z}\|$ is used. The proof is concluded.
VII-D Proof of Proposition 2
We construct the state feedback $v=\alpha_{v}(\bar{\zeta})$, with $\bar{\zeta}:=(\zeta,x_{11})$, which locally asymptotically stabilizes $\Sigma_{\bar{\zeta}}$ (14), at $\bar{\zeta}_{e}:=(\zeta_{e},x_{11e})$. Thus, from Theorem 10.3.1 in  [22], the cascaded interconnection of the autonomous system $\Sigma_{\bar{\zeta}}(\bar{\zeta};\alpha_{v}(\bar{\zeta}))$ and $\Sigma_{\eta}(\eta;\bar{\zeta})$ is guaranteed to be LAS in $\mathbb{U}_{\tilde{x}_{e}}\times SE(3)$ of $(\tilde{x}_{e},x_{11e})$, given the result in Theorem 1.
Consider the system $\Sigma^{H}_{3}(C^{0}_{1};\zeta_{1})$(75). The goal is to find a feedback law for the input $\zeta_{1}$ to stabilize the equilibrium $C^{0}_{1e}$. To this end, we choose the positive definite smooth Lyapunov candidate $V_{1}(C^{0}_{1}):SE(3)\rightarrow\mathbb{R}_{>0}\,\forall C^{0}_{1}\neq C^{0}_{1e},\,V_{1}(C^{0}_{1e})=0$ to be
[…]
$=\frac{1}{2}\operatorname{tr}\big((C^{1e}_{0}C^{0}_{1}(t)-I)^{T}(C^{1e}_{0}C^{0}_{1}(t)-I)\big)$
where $||.||_{F}$ is the Frobenius norm and $\operatorname{tr}(.)$ denotes the trace. The term $C^{1e}_{0}C^{0}_{1}(t)$ is the left-trivialized error on $SE(3)$. Define the variable $E_{c}$ as
$E_{c}=C^{1e}_{0}C^{0}_{1}(t)-I$ (86)
Taking the time derivatives of the $V_{1}(C^{0}_{1})$ along the trajectories of $C^{0}_{1}$ while simplifying using properties of the trace yields
$\dot{V}_{1}(C^{0}_{1}(t))$ $=\frac{1}{2}\frac{d}{dt}\operatorname{tr}\big(E_{c}^{T}E_{c}\big)=\frac{1}{2}\operatorname{tr}\big(\frac{d}{dt}E_{c}^{T}E_{c}+E_{c}^{T}\frac{d}{dt}E_{c}\big)$
$=\frac{1}{2}\operatorname{tr}\big(\frac{d}{dt}E_{c}^{T}E_{c}\big)+\frac{1}{2}\operatorname{tr}\big(E_{c}^{T}\frac{d}{dt}E_{c}\big)$
$=\frac{1}{2}\operatorname{tr}\big(E_{c}^{T}\frac{d}{dt}E_{c}\big)^{T}+\frac{1}{2}\operatorname{tr}\big(E_{c}^{T}\frac{d}{dt}E_{c}\big)$
$=\operatorname{tr}\big(E_{c}^{T}\frac{d}{dt}E_{c}\big)$ (87)
Since $\frac{d}{dt}E_{c}$ is given by
$\dot{E}_{c}=C^{1e}_{0}\frac{d}{dt}C^{0}_{1}(t)=C^{1e}_{0}[\zeta_{1}]{C}^{0}_{1}(t)$ (88)
Hence
$\dot{V}_{1}(C^{0}_{1}(t))$ $=\operatorname{tr}\big(E_{c}^{T}C^{1e}_{0}[\zeta_{1}]{C}^{0}_{1}(t)\big)=\operatorname{tr}\big({C}^{0}_{1}(t)E_{c}^{T}C^{1e}_{0}[\zeta_{1}]\big)$
$=\operatorname{tr}\big(((C^{1e}_{0})^{T}E_{c}({C}^{0}_{1})^{T})^{T}[\zeta_{1}]\big)$ (89)
By using the fact that if $b\in se(3),\,A\in\mathbb{R}^{4\times 4}$ then this trace property holds $\operatorname{tr}\big(A^{T}\,[b]\big)=\operatorname{tr}\big(P^{T}_{se(3)}(A)\,[b]\big)$ where $P_{se(3)}(A)$ projects the matrix $A$ on $se(3)$ as follows
$P_{se(3)}(A)=\left(\begin{array}[]{cc}\frac{1}{2}(A_{[1:3,1:3]}-A^{T}_{[1:3,1:3]})&A_{[1:3,4]}\\ 0_{1\times 3}&0\end{array}\right)$
where $A_{[a:b,c,d]}$ constructs a new matrix by extracting from matrix A the rows number $a$ to $b$ and columns number $c$ to $d$, $\dot{V}_{1}$ hence becomes
$\begin{split}\dot{V}_{1}(C^{0}_{1}(t))&=\operatorname{tr}\big(P^{T}_{se(3)}((C^{1e}_{0})^{T}E_{c}({C}^{0}_{1})^{T})[\zeta_{1}]\big)\\ &=\hat{P}^{T}_{se(3)}((C^{1e}_{0})^{T}E_{c}({C}^{0}_{1})^{T})L\zeta_{1}\end{split}$ (91)
where $L=\operatorname{diag}(2,2,2,1,1,1)$. Now let the input $\zeta_{1}$ be
$\zeta_{1}=-K_{p}\hat{P}_{se(3)}\Big((C^{1e}_{0})^{T}E_{c}({C}^{0}_{1})^{T}\Big)=:\zeta_{1e}$ (92)
where the diagonal gain matrix $K_{p}>0\in\mathbb{R}^{6\times 6}$ and $\hat{.}$ denotes the inverse map between $se(3)$ $\rightarrow$ $\mathbb{R}^{6}$. Substituting (92) makes $\dot{V}$ take this quadratic form
[…]
Now, we follow a backstepping strategy to stabilize the $\zeta_{1}$-dynamics at the equilibrium $\zeta_{1e}$ (92) using the input $\zeta_{2}$. Consider the positive definite smooth Lyapunov candidate $V_{2}({C}^{0}_{1},\zeta_{1}):SE(3)\times se(3)\rightarrow\mathbb{R}_{>0},\,\forall C^{0}_{1}\neq C^{0}_{1e}$ and $\forall\zeta_{1}\neq\zeta_{1e},\,\text{ with }V(C^{0}_{1e},\zeta_{1e})=0$
[…]
Taking the time derivatives of $V_{2}$ gives
$\begin{split}\dot{V}_{2}({C}^{0}_{1},\zeta_{1})&=\dot{V}_{1}+(\dot{\zeta}_{1}-\dot{\zeta}_{1e})^{T}(\zeta_{1}-\zeta_{1e})\end{split}$ (95)
hence by substituting the dynamics of $\zeta_{1}$ (74), i.e., $\dot{\zeta}_{1}=\zeta_{2}$
$\begin{split}\dot{V}_{2}({C}^{0}_{1},\zeta_{1})&=\dot{V}_{1}+(\zeta_{2}-\dot{\zeta}_{1e})^{T}(\zeta_{1}-\zeta_{1e})\end{split}$ (96)
and selecting the input $\zeta_{2}$ as
$\zeta_{2}=\dot{\zeta}_{1e}+\bar{\zeta}_{2}$ (97)
thus the derivatives of $V_{2}$ reads as
$\begin{split}\dot{V}_{2}({C}^{0}_{1},\zeta_{1})&=\dot{V}_{1}+\bar{\zeta}_{2}^{T}(\zeta_{1}-\zeta_{1e})\end{split}$ (98)
Setting $\bar{\zeta}_{2}$ to be $\bar{\zeta}_{2}=-K_{d}({\zeta}_{1}-\zeta_{1e})$, where the diagonal gain matrix $K_{d}>0\in\mathbb{R}^{6\times 6}$, renders $\dot{V}_{2}$ negative definite $\forall$ $(C^{0}_{1},\zeta_{1})\neq(C^{0}_{1e},\zeta_{1e})$
$\begin{split}\dot{V}_{2}({C}^{0}_{1},\zeta_{1})&=\dot{V}_{1}-({\zeta}_{1}-\zeta_{1e})^{T}K_{d}(\zeta_{1}-\zeta_{1e})<0\end{split}$ (99)
The total control law for ${\zeta}_{2}$ is
[…]
This step is repeated once again to design a controller $v$ that stabilizes the $\zeta_{2}$-dynamics (74) at the equilibrium $\zeta_{2e}$ with the Lyapunov candidate $V_{3}({C}^{0}_{1},\zeta_{1},\zeta_{2})$ given by
[…]
The equation for the virtual input $v$ is thus obtained as
[…]
where $K_{a}$ is a diagonal gain matrix $K_{a}>0\in\mathbb{R}^{6\times 6}$.
The time derivatives of ${\zeta}_{1e}$ are functions of the first time derivative of the error $E_{c}$ (86), i.e. $\dot{E}_{c}$ (88), and
$\ddot{E}_{c}$ $=C^{1e}_{0}\Big([\zeta_{2}]+[\zeta_{1}]^{2}\Big){C}^{0}_{1}$ (103)
Hence
$\begin{split}\dot{\zeta}_{1e}&=-K_{p}\hat{P}_{se(3)}\Big((C^{1e}_{0})^{T}\Big(\dot{E}_{c}({C}^{0}_{1})^{T}+E_{c}({C}^{0}_{1})^{T}[\zeta_{1}]^{T}\Big)\Big)\\ \ddot{\zeta}_{1e}&=-K_{p}\hat{P}_{se(3)}\Big((C^{1e}_{0})^{T}\Big(\ddot{E}_{c}({C}^{0}_{1})^{T}+2\dot{E}_{c}({C}^{0}_{1})^{T}[\zeta_{1}]^{T}+E_{c}({C}^{0}_{1})^{T}(([\zeta_{1}]^{T})^{2}+[\zeta_{2}]^{T})\Big)\Big)\end{split}$
After some algebraic manipulations of (100) and (102), the virtual control $v$ can be rewritten as
$v=\ddot{\zeta}_{1e}+\bar{K}_{d}(\dot{\zeta}_{1e}-{\zeta}_{2})+\bar{K}_{a}({\zeta}_{1e}-{\zeta}_{1})$ (105)
where $\bar{K}_{a}>0\in\mathbb{R}^{6\times 6},\,\bar{K}_{d}>0\in\mathbb{R}^{6\times 6}$ are diagonal gain matrices. They are related to ${K}_{a}$ and ${K}_{d}$ by
$\bar{K}_{a}={K}_{d}\,{K}_{a},\,\,\bar{K}_{d}={K}_{d}+{K}_{a}$ (106)
If the error states are defined as
$e_{1}(t):={\zeta}_{1e}-{\zeta}_{1},\,\,e_{2}(t):=\dot{e}_{1}$ (107)
then a linear error system with a state $e=(\,\,e_{1}\,\,e_{2}\,\,)^{T}$ and a pair $(A,B)$ given in Proposition 2 results. The gain matrices $\bar{K}_{d}$ and $\bar{K}_{a}$ can thus be tuned to exponentially stabilizes its origin by rendering (ABK)(A-BK) Hurwitz.
Remark 3.
Observe that to eliminate the steady-state error in the convergence of $e(t)$ to zero, which may be present due to parametric uncertainty, an integral term can be added to $v$ as
$\begin{split}v=\ddot{\zeta}_{1e}+\bar{K}_{d}(\dot{\zeta}_{1e}-{\zeta}_{2})+\bar{K}_{a}({\zeta}_{1e}-{\zeta}_{1})+K_{I}\int({\zeta}_{1e}-{\zeta}_{1})\,dt\end{split}$
where ${K}_{I}={K}^{T}_{I}>0\in\mathbb{R}^{6\times 6}$ is the integral gain matrix.
This concludes the proof.