### v3

# The PenduMAV: A Six-Input Omnidirectional MAV without Internal Forces - Design, Dynamics, and SE(3) Control

Abstract
We introduce the PenduMAV, an exactly actuated (6-input) omnidirectional multirotor that structurally eliminates internal forces at equilibria. The vehicle features one actively tilting propeller and three propellers mounted on passive pendulum links via universal joints. This architecture achieves full 6D wrench generation while avoiding the structural and energetic costs of input redundancy and internal forces. After deriving the full multibody dynamics, we demonstrate that a forced equilibrium exists for every main platform pose. To asymptotically stabilize the closed-loop system, we design a coordinate-invariant nonlinear controller based on dynamic feedback linearization and backstepping, utilizing the left-trivialized error on $SE(3)$. System stability is formally guaranteed through Lyapunov analysis of the zero dynamics. Finally, Gazebo simulations (videos available at https://www.youtube.com/playlist?list=PL4N8pJgvqASQX6AWEpg3NCZ6QdGBPfbXq) validate the approach, showcasing fully decoupled attitude and translational tracking under parametric uncertainty and actuator noise.
I introduction
Nowadays, Multirotor Aerial Vehicles (MAVs) are widely used in a variety of applications, ranging from agriculture to civil engineering  [1]. Interest in this technology is growing, geared by its potential to perform tasks that require physical interaction with environments inaccessible to other types of mobile vehicles [2, 3]. While conventional quadrotors can efficiently handle some of these tasks, particularly when special orientation maneuvers are not required, certain applications have emerged in which the capability of executing decoupled translational and rotational motions is essential, necessitating the demand for new MAV designs since quadrotor dynamics inherently prohibits this kind of motion. One notable example where the MAV platform needs to maintain constant orientation while translating in space is the contact-based inspection of sloped surfaces  [4, 5].
Research Problems and Related Work
When designing a new MAV, one can easily be convinced that it is desirable to obtain all the angular velocities of the propellers aligned with each other and with the gravity force at any static hovering equilibrium: this avoids internal propeller-generated forces canceling each other at the cost of a higher energy consumption; moreover, it is desirable to keep the number of actuators to the minimum (equal to the task dimension), to reduce the weight (which, in turn, affects the vehicle’s endurance), the cost, the maintenance requirements, etc. Furthermore, for tasks requiring high dexterity like the ones mentioned before, it is desirable for the MAV to be able to statically hover at any pose (in SE(3) in general). Last but not least, the design of a MAV must also allow the controlled stabilization of its hovering equilibria. We label these four desirable properties as the Internal-force Property, the Input-dimension Property, the Equilibria Property, and the Stabilizability Property, respectively.
When restricting the problem to the vertical 2D space, we demonstrated in [6] that a MAV design with all the above four properties exists. Especially, it can be stabilized at any static equilibrium pose in SE(2) with propellers parallel to each other and the gravity vector, and it has only three actuators. The actuators are two brushless motors attached to one propeller each and one servomotor that actively tilts one of the two propellers, while the other one is free to passively rotate and has a center of mass not coincident with the rotation point. However, extending such a result to the 3D space is very far from being trivial and is an open problem. Just to mention two obstacles to such extension: SE(3) is a fundamentally different manifold than SE(2) and the dynamic effects related to single revolute joints in 2D cannot be extended to the case of multiple joints in 3D, see model in Sec. II-C. Moreover, in the 3D space, the propeller drag torques arise and have to be taken into account, see discussion in Sec. III-A.
State of the Art
We revise the main MAV designs proposed in the literature and show that none of them has all the four desirable properties described above.
Standard underactuated MAVs like quadrotors and hexarotors tick the Internal-force Property as they have all propellers fixedly oriented with spinning axes parallel to each other; however, because of the fixed attachment between the propellers and the body frame, the only hovering equilibria are characterized by the attitudes where the $z$ axis is aligned with gravity. Thus, they do not tick the Equilibria property.
MAVs equipped with the ability to hover while maintaining a larger set of orientations, known as fully-actuated MAVs, are surveyed in [7]. These MAVs cannot be stabilized at any pose but at a subset of poses in SE(3) determined by the thrust limits, the design kinematics, and joint limits. Such subset contains any position but only a subset of all the orientations in general. A subclass of this type of MAVs, called omnidirectional MAV [8], can hover at any pose: namely, they have the Equilibria property mentioned above. These aerial vehicles can be divided in two classes.
The first class of designs that have been proposed to achieve full actuation or omnidirectionality are realized by fixedly tilting the propellers in such a way that their spinning axes are not all parallel to each other. As a consequence, all of these designs do not have the Internal-force Property. Examples are [9], a fully actuated hexarotor, and [10, 11], omnidirectional hexarotors. Other omnidirectional fixedly-tilted-propeller designs do not comply with the Input-dimension Property, either; they are, for example, the hepta-rotor of [12] or the octarotors proposed in [13, 14].
The second design approach to achieve full-actuation or omnidirectionality while also partially achieving the Internal-force Property for at least some hovering orientations is the addition of servomotors that actively change the relative orientation between the propellers and the frame of the MAV. In [15, 16], the authors present a modification of the conventional quadrotor by actuating the tilt of each propeller separately via 4 servomotors for a total of 8 actuators. Similarly, in [17], the authors used 6 servomotors in addition to the 6 brushless motors that rotate the 6 propellers of an hexarotor. The design in [18] also falls in this category. They developed a MAV with multiple links connected in series by active revolute joints, with each link having two propellers that rotate simultaneously using a motor. In an attempt to limit the number of actuators, some designs have been proposed in which a single servo-motor is used to synchronously tilt all the propellers: 6 propellers are synchronously tilted by the additional actuator in [19], and 8 propellers in [20]. All the aforementioned designs do not comply with the Input-dimension Property, having a number of actuators greater than 6. Furthermore they do not comply with the Internal-force Property at any orientation but only some at some specific ones.
Contributions
The main contribution of this work is to propose the PenduMAV concept, the first MAV design, to the best of the authors’ knowledge, having the following four properties:
  1. 1.
    Input-dimension property. The number of actuators in the proposed MAV design is 6, equal to the number of DoFs of its base’s pose. There are 4 brushless motors, each making one propeller spin; 3 of the propellers are connected to a passive 2-DoF joint; one of the propeller’s attitude is actively controlled by 2 servomotors.
  2. 2.
    Equilibria property. In Proposition 1, we show that static hovering is feasible at any pose in $SE(3)$ of the MAV’s main body; in other words, a forced equilibrium exists for any base’s pose.
  3. 3.
    Internal-force property. Proposition 1 also shows that no internal forces are produced at the equilibrium for any base pose in $SE(3)$.
  4. 4.
    Stabilizability property. We prove in Theorem 1 that it is possible to create a state feedback law that renders the closed-loop system of the MAV asymptotically stable at any base pose in $SE(3)$, as long as configuration and thrust trajectories of the system remain in a certain neighborhood of that base pose. We observe that the dynamics of the passive propeller joints coincide with the zero dynamics, and we prove its asymptotical stability using the Lyapunov theory. The feedback control law is given in Proposition 2.
Eventually, simulation results of the closed-loop system in the presence of non-idealities, such as noise and parameter uncertainties, are presented, in order to provide an acceptable level of confidence in the generalizability of the result beyond the standard model used to formally prove all the properties.
These properties capture major requirements for enabling energy-efficient omnidirectional flight. In particular, the ability to hover at any attitude in $SO(3)$ is the defining feature of omnidirectional MAVs [14, 15], hence expanding their reachable workspace and manipulation capabilities [12]. Ensuring stabilizability of these equilibria is essential for any practical deployment. Moreover, achieving full 6-DOF control with only 6 actuators alongside eliminating internal forces at hovering directly contributes to significantly improved energy efficiency and extended flight endurance [21, 13, 22]. Together, these properties illustrate the impact of the proposed MAV concept within the field of aerial robotics.
The rest of the paper is organized as follows. Sec. II describes the dynamic model of the proposed novel MAV conceptual idea, whereas Sec. III identifies the equilibria sets of that model being all the poses in SE(3) and all having the internal force property. Afterwards, we investigate and demonstrate that the system can be made asymptotically stable at all these equilibria by a state feedback in Sec. IV. Numerical simulations where the MAV performs decoupled motions are the subject of Sec. V. We conclude by giving some remarks on the future work in Sec. VI. For a streamlined presentation, the formal proofs, derivations, and explicit calculation formulas are collected in the appendices of [29].
TABLE I: Nomenclature
Symbol Description
$SE(3)$ Special Euclidean group of rigid-body transformations.
se(3){se}(3) Lie algebra of $SE(3)$.
$\mathbb{T}^{n}$ $n$-torus manifold (a compact Lie group).
$C^{0}_{j}\in SE(3)$ Configuration of $B_{j}$’s frame $\mathcal{F}_{p_{j}}$ in the world frame $\mathcal{F}_{w}$.
$V^{0}_{1}\in\mathbb{R}^{6}$ Right-invariant base twist expressed in, and relative to, $\mathcal{F}_{w}$.
[][*] $\mathbb{R}^{6}\to se(3)$ An isomorphism with an inverse $\hat{*}$.
$(*)_{\times}$ The skew-symmetric matrix of the vector $\in\mathbb{R}^{3}$.
$J_{1j}$ Universal joint connecting the base to link $B_{j}$.
$S_{j}\in\mathbb{R}^{6\times 2}$ $J_{1j}$’s spatial instantaneous screw axis, $S_{j}=(\,S_{j1}\;S_{j2}\,)$.
$S_{p}$ Block-diagonal matrix of screws $S_{j}$, $\operatorname{blkdiag}(S_{j})$.
$\operatorname{Ad}_{C}$ $SE(3)\times se(3)\rightarrow se(3)$ Adjoint representation of $SE(3)$.
$\operatorname{ad}_{V}$ $se(3)\times se(3)\rightarrow se(3)$ Adjoint representation of $se(3)$.
$R_{k}(*)$ Pure rotation about axis $k$ by a angle *.
$e_{i}$ Zero vector with a one at index $i$.
$M$ Generalized mass (inertia) matrix.
$M_{bb},M_{bp},M_{pp}$ Base, coupling, and joint sub-matrices of $M$.
$h$ Centrifugal and Coriolis terms.
$g_{b},g_{p}$ Gravitational terms of base and joints, $g=(\,g_{b}\,\,g_{p}\,)^{T}$.
$W^{0}_{\mathrm{app,b}}$ Applied wrench on base’s CoM expressed in $\mathcal{F}_{w}$.
$W^{0}_{\mathrm{app,p}}$ Applied wrench on links $B_{2},B_{3},B_{4}$’s CoM in $\mathcal{F}_{w}$ projected onto respective joints via $S_{p}$.
$\tau_{dr_{i}}$ Drag torque about propeller $i$’s axis.
$\tau_{f_{ji}}$ Viscous friction at joint $q_{ji}$’s axis. $\bar{\tau}_{f}=\big(0_{6\times 1}\;\;\operatorname{col}(\tau_{f_{ji}})\big)^{T}$.
$\tau_{s_{j}}$ Torsion spring torque applied at joint $q_{j1}$’s axis.
$\tau_{dj}$ Propeller $j$’s drag torque factorized on $q_{j1}$’s axis at equilibrium.
$m_{b},m_{j}$ Mass of base and link $B_{j}$.
$c$ Length from joints axes intersection point to $B_{j}$’s CoM.
$k_{i}$ Drag-to-thrust ratio of propeller $i$.
$\mathrm{PJD}_{1},\mathrm{PJD}_{2}$ Passive-joint designs of type 1 and type 2.
$P_{se(3)}(A)$ Projects a matrix $A\in\mathbb{R}^{4\times 4}$ onto $se(3)$.
$_{e}$ Variable * evaluated at equilibrium.
$^{[i]}$ Value of variable * pertaining to joint design PJDi.
II Proposed Design and Dynamical Model
In this section, we describe the proposed design and present its dynamical model in state-space form.
II-A Motivations Behind the Design Idea
The simplest MAV design that one could think of that respects the Input-dimension Property and the Internal-force Property (and potentially the Equilibria Property) is the one with 6 propellers, all free to passively tilt like 3D pendulums. However, the results achieved for the 2D example in [6] showed that a design with 3 passively linked propellers in a pendulum like fashion, in that case, was non-feedback linearizable as the allocation matrix was structurally rank-deficient at the static equilibria. Based on such a preliminary result, we discarded the all-passive-propeller design.
Hence, our design has servomotors that actively reorient one propeller. An alternative could be to reorient actively all the four propellers synchronously by means of the two servomotors and a transmission mechanism. However, the design of such a mechanism to drive the whole propeller joints synchronously is not trivial, would bring mechanical complexity and with it additional payload, cost, and maintenance effort; furthermore, it would not scale well with the size of the vehicle, when the motion needs to be transferred over longer distances. Instead, the proposed design where the servomotors exert their actuation locally at one individual propeller has the advantage of being modular and simple from a mechanical point of view, not posing challenges when increasing the dimensions of the vehicle.
II-B Mechanical Design Idea
Figure 1: Conceptual representation of the PenduMAV idea. The origins of the frames $\mathcal{F}_{b}$ and $\mathcal{F}_{p_{2}}$ are attached to CoMs of the base $B_{1}$ and the propeller link $B_{2}$, respectively. The shown configuration is when $\mathcal{F}_{b}$ coincides with $\mathcal{F}_{w}$ and all the frames of the links have the same orientation as $\mathcal{F}_{w}$. The joint $q_{j1}$ and $q_{j2}$ positions are zero when $\mathcal{F}_{p_{j}}$ has the same orientation as $\mathcal{F}_{b}$, $j=\{2,\ldots,5\}$. The link $B_{5}$ is attached to $B_{1}$ through an active joint, unlike the the rest of the joints which are passive.
Figure 2: The connection between $B_{j}$ ($j=2,3,4$) and the base $B_{1}$ in PJD2 consists of universal passive joint with a Pulley-Belt-Spring module. $\bar{B}_{j}$ is the intermediate link connecting $B_{1}$ to $B_{j}$. $b$, $p$ and $s$ denote the belt, pulley and a torsion spring, respectively. $\tau_{sj}$ is the torsion spring torque around the rotation axis of $q_{j1}$, i.e. $e_{j1}$.
The PenduMAV concept, shown in Fig. 1, consists of a main rigid body, referred to as the base and denoted by $B_{1}$, and four smaller bodies, referred to as links and denoted by $B_{2},B_{3},B_{4}$, and $B_{5}$. Each link carries a rigidly-attached motor driving a propeller and is connected to the base by two revolute joints. The propeller, the motor, the link, and the two revolute joints together form an actuation unit. For each unit, the joints’ mutually-orthogonal rotation axes intersect at a point, called the center of rotation of the link, coinciding with the propeller center. The actuation units are mounted on $B_{1}$ such that all centers of rotation lie on the same plane, resulting in a coplanar design. The centers of mass (CoMs) of $B_{2},B_{3}$, and $B_{4}$ are located below their respective centers of rotation, achieved by placing the motors sufficiently low. This makes these links behave as spatial pendula, with the joints acting as pivots. In contrast, $B_{5}$’s CoM coincides with its center of rotation. $B_{1}$ is connected to $B_{5}$ by an actuated universal joint with configuration $q_{a}=\operatorname{col}(q_{5i})\in\mathbb{T}^{2}$, and to $B_{2},B_{3}$, and $B_{4}$ by passive universal joints with configurations $q_{p}=\operatorname{col}(q_{ji})\in\mathbb{T}^{6}$, $j=2,\ldots,5$, $i=1,2$.
For modeling simplicity, the actuated joint is assumed to be driven by fast velocity-controlled servo DC motors, neglecting motor dynamics. Moreover, $B_{5}$’s mass is assumed to be concentrated at its CoM, zeroing its angular inertia moment. Under these assumptions, the projected dynamics of $B_{5}$} onto its $q_{a}$’s screw axis vanish, rendering $\dot{q}_{a}$ governed solely by the motor input.
Two Passive Joint Designs (PJD) are considered. In PJD1, the passive joints are implemented as described above. In PJD2, an additional Pulley-Belt-Spring module (Fig. 2) is integrated into each passive joint. This module transfers the rotation about $q_{j2}$’s axis to a rotation about $q_{j1}$’s axis through a belt-pulley transmission. A torsion spring, mounted along $q_{j1}$’s axis, applies a moment proportional to $q_{j2}$, thus the mechanism is unidirectional: rotations about $q_{j2}$ generate a restoring moment about $q_{j1}$, while the reverse is not true. The rationale behind PJD2 is discussed in Sec. III-A. Unless otherwise stated, the analysis herein applies to both designs.
II-C Equations of Motion
We denote with $\mathcal{Q}=(C_{1}^{0},{\color[rgb]{0,0,0}{q_{p}}},{\color[rgb]{0,0,0}{q_{a}}})\in SE(3)\times\mathbb{T}^{8}$ the system’s configuration. The velocity of $\color[rgb]{0,0,0}{(C_{1}^{0},q_{p})}$} is ${\color[rgb]{0,0,0}{(V_{1}^{0},\dot{q}_{p})}}\in\mathbb{R}^{12}$ where $V_{1}^{0}]=\dot{C}^{0}_{1}(C^{0}_{1})^{-1}\in se(3)$. Denote the full state vector $x(t)$ as $x=({\color[rgb]{0,0,0}{C_{1}^{0},q_{p}}},{\color[rgb]{0,0,0}{V_{1}^{0},\dot{q}_{p}}},{\color[rgb]{0,0,0}{q_{a}}})^{T}\in\mathbb{X}\equiv SE(3)\times\mathbb{T}^{8}\times\mathbb{R}^{12}$. Dropping dependencies for clarity, the system dynamics is thus described by these state space equations
[…]
where explicit formulas of $M,h$ and $g$ are given in Sec. -A [29], and $D_{b}$, $\color[rgb]{0,0,0}{D_{p}}$} and the input vector $u$ are defined as
$\begin{split}D_{b}=&\left(\begin{array}[]{cccc}\operatorname{Ad}^{T}_{(C^{0}_{f_{2}})^{-1}}\bar{f}_{2}&\operatorname{Ad}^{T}_{(C^{0}_{f_{3}})^{-1}}\bar{f}_{3}&\operatorname{Ad}^{T}_{(C^{0}_{f_{4}})^{-1}}\bar{f}_{4}&\operatorname{Ad}^{T}_{(C^{0}_{f_{5}})^{-1}}\bar{f}_{5}\end{array}\right)_{6\times 4}\\ {\color[rgb]{0,0,0}{D_{p}}}=&\left(\begin{array}[]{ccc}S_{2}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{2}})^{-1}}\bar{f}_{2}&0&0\\ 0&S_{3}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{3}})^{-1}}\bar{f}_{3}&0\\ 0&0&S_{4}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{4}})^{-1}}\bar{f}_{4}\end{array}\right)_{6\times 3}\\ u=&\left(\begin{array}[]{cccccc}f_{2}&f_{3}&f_{4}&f_{5}&u_{v_{1}}&u_{v_{2}}\end{array}\right)_{6\times 1}^{T}\in\mathbb{R}^{6}\end{split}$
[…]
Result 1 (Input dimension property).
The proposed design satisfies the input dimension property. In fact the dimension of the manifold of the configuration of the base frame $C_{1}^{0}$ is six and the dimension of the input is $\operatorname{dim}(u)=6$ (four propeller thrusts and two servomotor speeds).
III Study of Attainable Equilibria and
Internal Forces at each Equilibrium
In this section we obtain the set of forced equilibria of the dynamics (1) and investigate their properties in terms of the corresponding four thrust forces produced at these configurations. We begin by giving some preliminary definitions. Afterwards, the main result, which establishes the Equilibria property and the Internal-force property, is summarized in Proposition 1 whose proof is found in Sec. -D of [29].
Imposing $\dot{x}{(t)}=0$ in (1) implies $({\color[rgb]{0,0,0}{V_{1}^{0},\dot{q}_{p}}},{\color[rgb]{0,0,0}{\dot{q}_{a}}})=0$, which defines an equilibrium condition at which the state and the input become, respectively, $x_{e}=(C^{0}_{1e},{\color[rgb]{0,0,0}{q_{pe}}},0,0,{\color[rgb]{0,0,0}{q_{ae}}})^{T}$ and $u_{e}=(f_{e}\,\,u_{ve})^{T}$, with $f_{e}=(f_{2e}\,\,f_{3e}\,\,f_{4e}\,\,f_{5e})^{T}$ and $u_{ve}=(u_{v_{1e}}\,\,u_{v_{2e}})^{T}$. Under this condition, both $h({\color[rgb]{0,0,0}{C_{1}^{0},q_{p}}},{\color[rgb]{0,0,0}{V_{1}^{0},\dot{q}_{p}}})=0$ and $\bar{\tau}_{f}({\color[rgb]{0,0,0}{\dot{q}_{p}}})=0$ vanish, as they depend quadratically and linearly, respectively, on the configuration velocity.
For PJD1, (1) reduces to this static balance vector equation whose admissible solutions are the equilibrium state and input pairs $(x_{e},u_{e})$ such that
$g_{b}(C^{0}_{1e},{\color[rgb]{0,0,0}{q_{pe}}})$ $=(\,\,{\color[rgb]{0,0,0}{D_{b}}}\,\,0_{6\times 2}\,\,)u_{e},$ (3)
${\color[rgb]{0,0,0}{g_{p}}}(C^{0}_{1e},{\color[rgb]{0,0,0}{q_{pe}}})$ $=(\,\,{\color[rgb]{0,0,0}{D_{p}}}\,\,0_{6\times 3}\,\,)u_{e}$
$=(\,\,\tau_{d2}\,\,0\,\,\tau_{d3}\,\,0\,\,\tau_{d4}\,\,0\,\,)_{6\times 1}^{T}$
$0_{2}$ $=u_{ve}$ (4)
where $\tau_{dj}=k_{j}f_{je}\sin(q_{j2e})$, with $j=2,3,4$. $C^{0}_{1e}$ and ${\color[rgb]{0,0,0}{q_{pe}}}=\operatorname{col}(\,q_{je})$, with $q_{je}=\operatorname{col}(\,q_{jie})$ $\forall j=2,3,4,\forall i=1,2$.
III-A Effect of the Aerodynamic Drag Torque at the Equilibrium
The presence of the torque $\tau_{dj}$ in (4) leads to the important consequence that the value of the joints $\color[rgb]{0,0,0}{q_{pe}}$} at the equilibrium is not solely determined by the gravitational torque at equilibrium $\color[rgb]{0,0,0}{g_{pe}}$}, as it is the case for the 2D MAV presented in  [6], but also by a component of the equilibrium thrust $f_{e}$ at equilibrium.
This residual torque $\tau_{dj}$ can be eliminated at any equilibrium by the spring introduced in PJD2 whose proposed implementation in Fig. 2 enables a torsion spring to counteract residual torque and passively return the link to its vertical configuration, zeroing the internal force as stated in proposition 1. In fact, we can tune each variable-stiffness torsion spring such that its stiffness $k_{s}=\frac{\partial\tau_{sj}}{\partial q_{j2}}$ varies linearly with $\cos(q_{j2})$, with a constant slope $\bar{k}=k_{j}f_{je}$, i.e. $k_{s}=\bar{k}\cos(q_{j2})$, which is feasible since $f_{je}$ is found to be constant for any equilibrium of PJD2 as long as the vehicle weight remains constant during flight as shown in Sec. -D of [29]. Hence, the spring needs not to change or be re-tuned between flights. As a result, this component $\tau_{dj}$ is canceled at any PJD2 equilibrium, i.e. when $\tau_{dj}=-\tau_{sje}$, transforming (4) to
${\color[rgb]{0,0,0}{g_{p}}}(C^{0}_{1e},{\color[rgb]{0,0,0}{q_{pe}}})=0_{6},\,\,\,u_{ve}=0_{2}.$ (5)
III-B Minimal Thrust at Equilibrium vs Presence of Internal Forces
Before presenting the result on the Equilibria and Internal-force properties of the proposed MAV (1) we provide two definition for the thrust magnitudes at the equilibrium, in the case in which they are minimal and in the case in which there is internal force.
Definition 1.
A vector of thrust magnitudes $f\in\mathbb{R}_{\geq 0}^{l}$, where $\mathbb{R}_{\geq 0}^{l}:=\left\{s=col(s_{i})\in\mathbb{R}^{l}\mid s_{i}\geq 0,\ \forall i\in\{1,\dots,l\}\right\}$, is said to be minimal at the equilibrium if it belongs to $\mathcal{F}_{\rm min}^{\rm eq}$, defined as:
$\mathcal{F}_{\rm min}^{\rm eq}:=\bigg\{f=col(f_{i})\in\mathbb{R}_{\geq 0}^{l}\bigg|\,\sum^{l}_{i=1}{f}_{i}=m_{\rm tot}\,g_{r}\bigg\}$ (6)
where $m_{\rm tot}$ is the robot’s mass and $g_{r}=9.81m/s^{2}$. Instead, $f$ is said to have internal forces at the equilibrium if it belongs to $\mathcal{F}_{\rm int.for.}^{\rm eq}$, defined as:
$\mathcal{F}_{\rm int.for.}^{\rm eq}:=\bigg\{f=col(f_{i})\in\mathbb{R}_{\geq 0}^{l}\bigg|\,\sum^{l}_{i=1}{f}_{i}>m_{\rm tot}\,g_{r}\bigg\}$ (7)
A MAV design whose thrust magnitudes are minimal at all its equilibria is said to have no internal forces at its equilibria.
In other words, a minimal thrust magnitudes vector $f_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ has the least amount of total thrust magnitudes necessary to counteract the gravity at the equilibrium, whereas $f_{e}\in\mathcal{F}_{\rm int.for.}^{\rm eq}$ means that the vehicle generates more total thrust than what is strictly necessary for compensating gravity, thus having some wasted internal force. Moreover, there exists a unique orientation for each propeller in an MAV such that when all the propellers assume such orientation the absence of internal force is guaranteed at hovering, proven in Lemma 1 in Sec. -C [29]. The following proposition establishes the main result of this section.
Proposition 1 (Equilibrium Set of the System).
Consider any pose for the base $B_{1}$ denoted by $C^{0}_{1e}$. Then the following statements hold true and are equivalent for the system (1):
  1. 1.
    For any base pose $C^{0}_{1e}\in SE(3)$, there exists an equilibrium for PJD2 in which the joint equilibrium configuration ${\color[rgb]{0,0,0}{q_{pe}}}^{[2]}$ satisfies $C^{0}_{j}(C^{0}_{1e},q^{[2]}_{je})e_{3}=e_{3}$ $\forall j\in\{2,\dots,5\}$ and a thrust magnitudes vector $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ is produced whose elements $f^{[2]}_{je}={g_{r}(\frac{1}{4}m_{b}+m_{j})}$ are constants $\forall C^{0}_{1e}$ with $k_{i}$ chosen by $\sum_{j=2}^{5}k_{i}{(\frac{1}{4}m_{b}+m_{j})}=0$.
  2. 2.
    For any base pose $C^{0}_{1e}$, PJD2 has no internal force at any equilibrium ${\color[rgb]{0,0,0}{q_{pe}}}^{[2]}$, by Definition 1, as a corollary of $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ (Statement 1).
  3. 3.
    When $\tau_{dj}\neq 0$, PJD1 has no joint equilibrium ${\color[rgb]{0,0,0}{q_{pe}}}^{[1]}$ in which each propeller produces a minimal thrust: $C^{0}_{j}(C^{0}_{1e},q^{[1]}_{je})e_{3}\neq e_{3},\,\forall\tau_{dj}\neq 0$.
  4. 4.
    When $\tau_{dj}\neq 0$, it follows from Statement 3, and Lemma 1 in [29], that PJD1 has an internal force $f^{[1]}_{e}\in\mathcal{F}_{\rm int.for.}^{\rm eq}$. The amount of such force can be reduced by changing parameters $k_{i}$,c\,c and $m_{j}$.
Proof.
See Sec. -D of [29]. ∎
The state equilibria set for PJDi is denoted by $\mathbb{D}^{[i]}_{x_{e}}$ whereas $\mathbb{D}^{[i]}_{f_{e}}$ signifies their associated thrust equilibrium sets. The explicit relations are given in Sec. -D4 of [29].
We conclude the section remarking that the perfect vertical alignment of propellers in any PJD2 equilibrium is possible thanks to each spring torque $\tau_{sje}$ absorbing the residual drag $\tau_{dj}$ affecting the joint variable $q_{j1}$. This allows any joint equilibrium ${\color[rgb]{0,0,0}{q_{pe}}}^{[2]}$ to be dictated only by the gravity acting on the CoMs of the pendulum-like links B2-B4, given that the pivots, i.e., the joints, are passive.
Remark 1.
The design using PJD1 constitutes still an excellent compromise between mechanical simplicity and energy efficiency, with the ability to outperform the state of the art designs when it comes to the amount of internal force needed at the the equilibria. In fact, for typical values of the aerodynamic and kinematic parameters, the difference in the amount of internal forces generated in PJD1 design remains significantly small compared to the ideal PJD2 case, as it can be seen for example in the simulations of Section V. This supports the viability of the concept from a mechanical perspective by using the simpler design PJD1, if one does not want to resort to the particular joint design in PJD2.
IV Stabilization of the Equilibria
Following the characterization of the Equilibria and Internal-Force properties in Proposition 1, this section presents the solution of the problem of whether every equilibrium point in the equilibria set $\mathbb{D}^{[i]}_{x_{e}}$ can be made closed-loop asymptotically stable, at least locally, by a regular state feedback law, hence establishing the Stabilizability property. The core findings are stated in Theorem 1 and Proposition 2. The proposed control architecture is shown in Fig. 3.
The analysis departs by noticing that there exists no static state feedback control law rendering the closed-loop system $\Sigma_{1}$ in (1) locally asymptotically stable (LAS) in an open neighborhood $\mathbb{U}_{{x}_{e}}\subseteq\mathbb{X}$ of an equilibrium ${x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}$, since it turns out that $\Sigma_{1}$ does not possess a well-defined vector relative degree at any $x_{e}\in\mathbb{D}^{[i]}_{x_{e}}$  [23]. However, the goal of stabilizing this closed-loop system at $x_{e}$ remains attainable by a dynamic state feedback obtained from the dynamic extension algorithm [24], as illustrated in Theorem 1.
Figure 3: Block diagram of the control scheme. It takes a desired base pose $C^{0}_{1e}\in SE(3)$ and converts it into a reference base twist $\zeta_{1e}(t)$ through configuration error feedback $E_{c}$. Dynamic FBL is applied to the twist dynamics, which are stabilized using a PID controller. The accelerations are efficiently computed via a modified Articulated-Body Inertia Algorithm (ABA) in Sec. -B of [29].
Theorem 1 (Feedback Linearization Compensator and Zero Dynamics (inner loop)).
Let the original system $\Sigma_{1}$ (1) be decomposed into $\Sigma_{2}$ and $\Sigma_{3}$, respectively given by $\Sigma_{3}:\dot{C}_{1}^{0}=[V_{1}^{0}]\,C_{1}^{0}$ and this $\Sigma_{2}$
$\Sigma_{2}:\left\{\begin{aligned} \dot{x}_{2}=&{\color[rgb]{0,0,0}{F_{2}(x)+G_{2}(x)u}}\\ =&\left(\begin{array}[]{c}{\color[rgb]{0,0,0}{\dot{q}_{p}}}\\ M^{-1}(-h-g+\bar{\tau}_{f})\\ 0_{2\times 1}\end{array}\right)+\left(\begin{array}[]{c}0_{6\times 6}\\ M^{-1}\left(\begin{array}[]{cc}D_{b}&0_{6\times 2}\\ D_{t}&0_{6\times 3}\\ \end{array}\right)\\ \begin{array}[]{cc}0_{2\times 4}&I_{2\times 2}\end{array}\end{array}\right)u\end{aligned}\right$.
Then, these statements are true:
  1. 1.
    $\Sigma_{2}$ is dynamically input/output feedback linearizable w.r.t the I/O pair $(u,h_{y})$, where the output function $h_{y}(t):t\rightarrow\mathbb{R}^{6}$ is $h_{y}(t)={\color[rgb]{0,0,0}{V_{1}^{0}}}$, by a dynamic extension [24] given by this input precompensator
    $z(t)=col(z_{i}):=f(t)\in\mathbb{R}^{4},\dot{z}=\bar{w}\in\mathbb{R}^{4},i=2,\ldots,5.$
    The extended state is $\bar{x}=({\color[rgb]{0,0,0}{q_{p}}},{\color[rgb]{0,0,0}{V_{1}^{0},\dot{q}_{p}}},{\color[rgb]{0,0,0}{q_{a}}},z)\in\bar{\mathbb{X}}\equiv\mathbb{T}^{8}\times\mathbb{R}^{16}$ whereas the linearizing state feedback on the extended system, whose I/O pair is $(w,h_{y})$, is given by
    ${w}=\left(\begin{smallmatrix}\bar{w}\\ u_{v_{1}}\\ u_{v_{2}}\end{smallmatrix}\right)=D_{2}^{-1}M_{b}(-E_{2}+v)$ (10)
    where $v\in\mathbb{R}^{6}$ is the virtual input, and $E_{2}$, $D_{2}$ and $M_{b}$ are given in (62), (55) and (60) of Sec. -E [29]. This dynamic FBL is well-defined in an open neighborhood $\mathbb{U}_{\bar{x}_{e}}\subseteq\bar{\mathbb{X}}$ of $\bar{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}\cup\mathbb{D}^{[i]}_{f_{e}}$ defined as
    $\mathbb{U}_{\bar{x}_{e}}=\Big\{\bar{x}(t)\in\bar{\mathbb{X}}\,\Bigr\rvert\,\operatorname{det}\big(D_{2}(\bar{x})\big)\neq 0\Big\}$ (11)
  2. 2.
    For any base pose ${\color[rgb]{0,0,0}{C_{1e}^{0}}}\in SE(3)$, The original system $\Sigma_{1}$, extended with the precompensator (9), has locally, in $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$, asymptotically stable (LAS) zero dynamics w.r.t the output $h_{y}$, iff $\forall\bar{x}\in\mathbb{Z}_{x}-\{\bar{x}_{e}\}$
    $\|\tilde{f}\|<\frac{1}{\|D_{3}({\color[rgb]{0,0,0}{q_{p}}}-{\color[rgb]{0,0,0}{q_{pe}}})\|}\lambda_{\text{min}}(D_{f})\|{\color[rgb]{0,0,0}{\dot{q}_{p}}}\|$ (12)
    where $\tilde{f}=col({f}_{i})\,,i=2,3,4$, $\mathbb{Z}_{x}$ and $D_{3}$ are respectively given in (71) and (65) in [29]. $D_{f}>0\in\mathbb{R}^{6\times 6}$ contains the constant passive joints friction coefficients.
Proof.
See Sec. -E in [29].∎
This means that, as long as the extended system trajectories remain in $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$ where the decoupling holds, any base pose $C^{0}_{1e}$ of the original system $\Sigma_{1}$ (1) can be made locally asymptotically stable by a dynamic state feedback. Proposition 2 constructs the the virtual controller that stabilizes the resulting linearized twist dynamics.
Proposition 2 (Asymptotic Stabilization of Equilibria (outer loop)).
Consider the virtual control input $v$ given by
$v=\ddot{\zeta}_{1e}+\bar{K}_{d}(\dot{\zeta}_{1e}-\dot{V}^{0}_{1})+\bar{K}_{a}({\zeta}_{1e}-V^{0}_{1})$ (13)
where $\zeta_{1e}$, $\dot{\zeta}_{1e}$, $\ddot{\zeta}_{1e}$ $\in\mathbb{R}^{6}$ are respectively given in (86) and (98[29] and shown in Fig. 3. $\bar{K}_{d}$ and $\bar{K}_{a}$ $\in\mathbb{R}^{6\times 6}$ are positive definite diagonal gain matrices chosen such that to make (ABK)(A-BK) Hurwitz where
$A=\left(\begin{array}[]{c}0_{6\times 6}\quad I_{6\times 6}\\ 0_{6\times 6}\quad 0_{6\times 6}\\ \end{array}\right),\,B=\left(\begin{array}[]{c}0_{6\times 6}\\ I_{6\times 6}\\ \end{array}\right),\,K=\left(\begin{array}[]{c}\bar{K}_{a}\quad\bar{K}_{d}\end{array}\right)$
Given the precompensator in (9) and the linearizing control law in (10), substituting $v$ into (10) asymptotically stabilizes the extended system $\Sigma_{1}$ with (9) in the open neighborhood $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$.
Proof.
See Sec. -F in [29]. ∎
We note that the system (1) evolves on the product Lie group $SE(3)\times\mathbb{T}^{8}$ due to the internal joint dynamics. This motivates our use of the normal-form approach [27], as it facilitates the study of the associated zero dynamics. Although the proposed controller, shown in Fig. 3, employs tools often considered coordinate-dependent, the Lie-group formulation guarantees that the control laws are coordinate-invariant, i.e., their validity does not depend on a specific local parameterization of the configuration $SE(3)\times\mathbb{T}^{8}$.
V Simulation Tests
This section reports high-fidelity Gazebo simulations, implemented via the GenoM3 framework, validating the proposed controller beyond the nominal model and under more realistic conditions such as presence of parametric uncertainty. For both designs PJD1 and PJD2, the controller (10), (13) with the dynamic extension (9), shown in Fig. 3, achieves independent stabilization of the MAV attitude and position. Two indicative desired base poses are considered: an attitude command of a $60^{\circ}$ step in roll, pitch and yaw at fixed position, and a $1$ [m] position step along the inertial $(x,y,z)$ with constant orientation. The resulting trajectories are shown in Fig. 4. Simulation videos and plots for several different maneuvers are available at https://www.youtube.com/playlist?list=PL4N8pJgvqASQX6AWEpg3NCZ6QdGBPfbXq. Most notably, a simulation with $270$-degree roll turn is included (see also Fig. 7and 8 [29]) which demonstrates full omnidirectional capabilities. As it can be noted from these simulations, the vehicle can achieve a wide range of maneuvers while generating only positive thrust, potentially reducing the complexity associated with the use of bidirectional propellers. The state and input evolutions exhibit the natural real-world-like oscillations due to the imperfections introduced by parametric uncertainty and actuator noise from the emulated hardware.
(a) PJD1
(b) PJD2
Figure 4: Evolutions of the state and input for (a) PJD1 and (b) PJD2 designs. The simulation has two phases: First (0-60s) is the position step with the initial attitude held constant. Second (60-120s, gray-shaded), the attitude step with the position at $t=60$s held constant.
[…]
To illustrate the stability of the zero dynamics (72[29], as proven in Sec. -E, we simulated its evolution under nominal conditions with initial values $C^{0}_{1}(0)=I_{4\times 4},\,{\color[rgb]{0,0,0}{q_{p}}}(0)=-30^{\circ}\times\boldsymbol{1}_{6\times 1}\,$, $\dot{{\color[rgb]{0,0,0}{q_{p}}}}(0)=0.1\times\boldsymbol{1}_{6\times 1}$ rad/s. As shown in Fig. 5, the $q_{p}$ and $\dot{q}_{p}$ converge to zero, confirming convergence to the equilibrium $q_{pe}$ corresponding to that base pose $C^{0}_{1e}(0)$.
Figure 5: zero dynamics evolution showing convergence.
Figure 6: Comparison of the total thrust between PJD1 and PJD2 when stabilizing a $60^{\circ}$ pitch attitude step. In hovering, $\sum f^{[1]}_{e}$ is 2% higher than $\sum f^{[2]}_{e}$.
[…]
VI Conclusion
Figure 7: Top: the PenduMAV, simulated with an independent and uncertain/noisy model in Gazebo with the DART physics engine, performing a 270-degree roll maneuver. Video with different maneuvers available at https://www.youtube.com/playlist?list=PL4N8pJgvqASQX6AWEpg3NCZ6QdGBPfbXq.
Bottom: Position and attitude time evolutions of the two phases.
This work proposes the PenduMAV a novel omnidirectional MAV concept enabling full 6-DoF motion with a minimum number of actuators and no internal forces at equilibrium. The design combines a rigid base evolving on $SE(3)$, thus avoiding representation singularities, with 4 propeller-carrying links connected through universal joints, 3 of which are passive whereas 1 is actively driven by 2 servomotors. We proved that the proposed MAV satisfies the Input-Dimension, Equilibria, Internal-Force, and Stabilizability properties. Without overactuation, an equilibrium exists for any base pose, and when torsion springs are employed, the total thrust exactly balances the vehicle weight. Even without springs, internal forces can be made arbitrarily small through changing the links masses and lengths. Any base pose can be rendered locally asymptotically stable via a dynamic FBL and backstepping from base pose error on $SE(3)$. High-fidelity simulations demonstrate decoupled position and attitude control under realistic conditions.
This design scales favorably, as hovering loads are distributed almost evenly among propellers, avoiding excessive stress on the actively tilting unit. Scaling challenges are therefore typical to those of multirotors, such as propeller and battery sizing, rather than limitations specific to the proposed MAV. Future works will focus on a large and long term project to realize, implement and test the mechanical concept.